### v1

# A Note on Sample Complexity of Interactive Imitation Learning with Log Loss

Abstract
Imitation learning (IL) is a general paradigm for learning from experts in sequential decision-making problems. Recent advancements in IL [4] have shown that offline imitation learning, specifically Behavior Cloning (BC) with log loss, is minimax optimal. Meanwhile, its interactive counterpart, DAgger, is shown to suffer from suboptimal sample complexity. In this note, we focus on realizable deterministic expert and revisit interactive imitation learning, particularly DAgger with log loss. We demonstrate: 1. A one-sample-per-round DAgger variant that outperforms BC in state-wise annotation. 2. Without recoverability assumption, DAgger with first-step mixture policies matches the performance of BC. Along the analysis, we introduce a new notion of decoupled Hellinger distance that separates state and action sequences, which can be of independent interest.
Yichen Li Chicheng Zhang
University of Arizona University of Arizona
yichenl@arizona.edu chichengz@cs.arizona.edu
1 Introduction
Imitation learning (IL) is a general paradigm for learning to make sequential decisions. Given offline demonstrations or an interactive expert, the goal is to learn policy that matches the performance of the expert. Two prominent frameworks in IL are offline IL [13] and interactive IL [15]. A popular method of offline IL, behavior cloning (BC), treats IL as a supervised learning problem by collecting trajectories from the expert and learning a policy that maps states to actions. However, it is well known that this may result in compounding error [12], where the learner makes an initial mistake and fails to recover form consecutive unseen states.
To address this challenge, interactive IL [15] introduces training-time expert interaction and utilizes it by iteratively collecting states visited by the learner and querying the expert for action annotations. Interactive IL has been previously known to outperform offline IL under the recoverability assumption [15, 14, 20, 9], where the expert can recover from action deviations with low extra cost.
Recent advancement [4] raises a critical question about the utility of interaction in imitation learning. Specifically, they show that offline IL achieves minimax-optimal sample complexity for both interactive and offline IL, where the sample complexity is in terms of the number of full trajectories annotated. The benefits of interaction for imitation learning highlighted in [4] are restricted to specific scenarios. In this note, we ask the following question:
Can interactive imitation learning be more sample-efficient than offline imitation learning in general?
To answer this, we revisit and conduct refined analysis of (variants) of existing interactive IL algorithms. First, we demonstrate that, when the expert provides state-wise annotations, under a low-recoverability setting, interactive IL can achieve significantly improved sample complexity compared to offline IL. Second, when the expert provides trajectory-wise annotations, we show that without the recoverability assumption, an interactive IL algorithm has sample complexity no worse than that of behavior cloning, contrary to prior claims [4]. Together, these results highlight the potential benefits of interaction in imitation learning and calls for a systematic understanding on imitation learning with state-wise annotations.
2 Preliminaries
Basic notation.
Define $[n]:=\left\{1,\ldots,n\right\}$. Denote by $\Delta(\mathcal{X})$ the set of probability distributions over a finite set $\mathcal{X}$. For $u\in\Delta(\mathcal{X})$ and $x\in\mathcal{X}$, we denote by $u(x)$ the $x$-th coordinate of $u$ and $e_{x}$ the delta mass on $x$. We use the shorthand $x_{1:H}$ to represent the sequence $\{x_{i}\}_{i=1}^{n}$.
Episodic Markov decision process and policy interaction.
[…]
Additional policy related notations.
[…]
Assumption 1 (Realizability).
The expert policy $\pi^{E}$ is contained in the learner’s policy class $\mathcal{B}$.
We will consider learning from two different policy classes induced by $\mathcal{B}$:
Definition 1 (First-step mixing of $\mathcal{B}$).
$\Pi_{\mathcal{B}}:=\left\{\pi_{u}:u\in\Delta(\mathcal{B})\right\}$, where policy $\pi_{u}$ is executed in an an episode of an MDP $\mathcal{M}$ by: draw $\pi\sim u$ at the beginning of the episode, and execute policy $\pi$ throughout the episode.
Importantly, $\pi_{u}$ is not a stationary policy; as a result, $a_{1:H}$ are dependent conditioned on $s_{1:H}$; $a_{1:H}$ are only conditionally independent given $s_{1:H}$ and the random policy $\pi$ drawn.
Definition 2 (Each-step mixing of $\mathcal{B}$).
$\bar{\Pi}_{\mathcal{B}}=\{\bar{\pi}_{u,h}(a|s):=\sum_{\pi\in\mathcal{B}}u(\pi)\pi_{h}(a|s):u\in\Delta(\mathcal{B})\}.$
In words, an each-step mixing policy $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$ can be executed by drawing $\pi\sim u$ freshly-at-random at each step $h\in[H]$ and plays $\pi_{h}(a|s)$ (e.g. [9, 10]). Observe that $\bar{\pi}_{u}$ is a Markovian policy.
Additionally, we use $\pi(\cdot\parallel s_{1:H})$ to denote the causally-conditioned probability of action sequence $a_{1:H}$ induced by $\pi$, given state sequence $s_{1:H}$ [21].  [^1] The use of $\parallel$ highlights its distinction from standard conditioning on $s_{1:H}$. To elaborate:
  • For Markovian policy $\pi$, $\pi(a_{1:H}\parallel s_{1:H}):=\prod_{h=1}^{H}\pi_{h}(a_{h}|s_{h})$.
  • For first-step mixing of Markovian policies $\pi_{u}$, $\pi_{u}(\cdot\parallel s_{1:H}):=\sum_{\pi\in\mathcal{B}}u(\pi)\pi(\cdot\parallel s_{1:H})$.
It is well-known that the trajectory distribution induced by Markovian policies and their first-step mixtures $\pi$ can be factorized to the product of $\pi(a_{1:H}\parallel s_{1:H})$ and the causally-conditioned probability of the state sequence given the action sequence (Definition 23 and Lemma 24). When it is clear from context, we use shorthand $\pi(s_{1:H})$ for $\pi(\cdot\parallel s_{1:H})$.
Interactive imitation learning protocol.
We consider two oracles that models interacting with demonstration expert: first, the trajectory-wise demonstration oracle $\mathcal{O}^{\mathrm{Traj}}$ that takes into a state-sequence $s_{1:H}$ and returns $a^{*}_{1:H}\sim\pi^{\mathrm{E}}(\cdot\parallel s_{1:H})$; second, the state-wise demonstration oracle $\mathcal{O}^{\mathrm{State}}$ that takes into a state $s$ and step $h$ and returns $a^{*}\sim\pi^{\mathrm{E}}_{h}(\cdot\parallel s)$.
We consider interactive IL with the learner interacting with the environment and demonstration oracles using the following protocol:
  • For $i=1,2,\dots$
    • Select policy $\pi^{i}$ and rollout in $\mathcal{M}$, observing trajectory $(s_{1},a_{1},\ldots,s_{H},a_{H})$.
    • Query the available oracle(s) to obtain expert annotations.
  • Goal: Return policy $\hat{\pi}$ such that $J(\pi^{\mathrm{E}})-J(\hat{\pi})$ is small, with a few number of queries to $\mathcal{O}^{\mathrm{Traj}}$ or $\mathcal{O}^{\mathrm{State}}$.
By replacing the query of $\mathcal{O}^{\mathrm{Traj}}$ on $s_{1:H}$ with $H$ separate queries to $\mathcal{O}^{\mathrm{State}}$, any algorithm that makes $N$ queries to $\mathcal{O}^{\mathrm{Traj}}$ can be (trivially) converted to one that makes $NH$ queries to $\mathcal{O}^{\mathrm{State}}$. We will later use this observation to make comparison in state-wise annotation model.
Our IL algorithm design will rely on basic online learning algorithms as primitives. An online learning oracle $\mathbb{A}$ with decision space $U$ is said to achieve a regret of $\operatorname{Reg}$ with benchmark set $V$, if it repeatedly chooses $u^{n}\in U$ and then observes loss $\ell^{n}:U\to\mathbb{R}$, and satisfies
$\sum_{n=1}^{N}\ell^{n}(u^{n})-\min_{v\in V}\sum_{n=1}^{N}\ell^{n}(v)\leq\operatorname{Reg}.$
3 Main Results
3.1 Interactive IL enables better sample complexity with state-wise annotation
We first present a interactive IL algorithm, Algorithm 1, that works by interacting with the state-wise annotation oracle $\mathcal{O}^{\mathrm{State}}$.
Algorithm 1 DAgger with state-wise annotation oracle
1:  Input: MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$ with decision space $\Delta(\mathcal{B})$ and benchmark set $\left\{e_{\pi}:\pi\in\mathcal{B}\right\}$.
2:  for $n=1,\dots,N$ do
3:     Query $\mathbb{A}$ and receive $u^{n}\in\Delta(\mathcal{B})$.
4:     Execute $\pi^{n}:=\bar{\pi}_{u^{n}}$ and sample $(s^{n},h^{n})\sim d^{\pi^{n}}$. Query $\mathcal{O}^{\mathrm{State}}$ for $a^{*,n}=\pi^{\mathrm{E}}_{h^{n}}(s^{n})$.
5:     Update $\mathbb{A}$ with loss function
$\ell^{n}(u):=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi_{h}(a^{*,n}|s^{n})}\right).$ (1)
6:  end for
7:  Output $\hat{\pi}$, a first-step uniform mixture of policies in $\left\{\pi^{1},\ldots,\pi^{n}\right\}$.
Algorithm 1 requires base policy class $\mathcal{B}$ and uses each-step mixing policies $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$ (recall Definition 2). At round $n$, it rolls out the current policy $\pi^{n}=\bar{\pi}_{u^{n}}$, whose mixing weight $u^{n}$ is obtained from an online learning oracle $\mathbb{A}$ and samples $(s^{n},h^{n})$ from $d^{\pi^{n}}$. The online learning oracle $\mathbb{A}$ it uses has decision space $\Delta(\mathcal{B})$ and benchmark set $\left\{e_{\pi}:\pi\in\mathcal{B}\right\}$; an classical example of this is the exponential weight algorithm, which has regret bound $\operatorname{Reg}=\ln B$ (see Proposition 21). It then queries $\mathcal{O}^{\mathrm{State}}$ to get $a^{*,n}$ and updates $\mathbb{A}$ by loss $\ell^{n}(u)$ induced by the new example (Eq. (1)). The final policy $\hat{\pi}$ is returned as a uniform first-step mixture of the historical policies $\{\pi^{n}\}_{n=1}^{N}$. We show the following performance guarantee of Algorithm 1:
Theorem 3.
If Algorithm 1 is run with a deterministic expert policy $\pi^{E}$, an MDP $\mathcal{M}$ such that $(\mathcal{M},\pi^{E})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that realizability holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm, then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq\mu H\cdot\frac{\log(B)+2\log(1/\delta)}{N}.$
Theorem 3 shows that Algorithm 1 returns a policy of suboptimality $O(\frac{\mu H\ln B}{N})$ using $N$ state-wise annotations from the expert. In comparison, with $N$ state-wise annotations, one can obtain $N/H$ trajectory-wise annotations; [4]’s analysis shows that behavior cloning on $N/H$ iid trajectories from $\pi^{\mathrm{E}}$ returns a policy of suboptimality $O(\frac{RH\ln B}{N})$. As $\mu$ is no greater than $R$ and can be much smaller, Algorithm 1 has a faster convergence rate than behavior cloning under the same state-wise annotation budget.
[…]
$\operatorname{OnEst}^{\mathrm{State}}_{N}:=\sum_{n=1}^{N}\mathbb{E}_{(s,h)\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}_{h}(s),\pi^{\mathrm{E}}_{h}(s))\right].$
The proof involves bounding the state-wise Hellinger estimation error and translating it into the performance difference difference beteween $\hat{\pi}$ and $\pi^{\mathrm{E}}$. More details are provided in Appendix C.
3.2 Interactive IL matches offline IL on trajectory-wise annotation
Next, we consider the trajectory-wise sampling model. We present another DAgger variant, namely Algorithm 2 and provide its sample complexity bounds.
Algorithm 2 DAgger with trajectory-wise annotation oracle
1:  Input: MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$ with decision space $\Delta(\mathcal{B})$ and benchmark set $\left\{e_{\pi}:\pi\in\mathcal{B}\right\}$.
2:  for $n=1,\dots,N$ do
3:     Query $\mathbb{A}$ and receive $u^{n}\in\Delta(\mathcal{B})$.
4:     Execute $\pi^{n}:=\pi_{u^{n}}$ and sample $s^{n}_{1:H}$ following $\mathbb{P}^{\pi^{n}}$. Query $\mathcal{O}^{\mathrm{Traj}}$ for $a_{1:H}^{*,n}=\pi^{\mathrm{E}}(s^{n}_{1:H})$.
5:     Update $\mathbb{A}$ with loss function
$\ell^{n}(u):=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi_{h}(a_{1:H}^{*,n}\parallel s_{1:H}^{n})}\right).$ (2)
6:  end for
7:  Output $\hat{\pi}$, the first-step uniform mixture of policies in $\left\{\pi^{1},\ldots,\pi^{n}\right\}$.
Algorithm 2 uses first-step mixing policies $\pi_{u}\in\Pi_{\mathcal{B}}$ (recall Definition 1). At round $n$, it rolls out $\pi^{n}=\pi_{u^{n}}$ whose mixing weight $u^{n}$ is obtained from an online learning oracle $\mathbb{A}$ and samples a full state sequence $s^{n}_{1:H}$. Same as Algorithm 1, Algorithm 2 also requires $\mathbb{A}$ to have decision space $\Delta(\mathcal{B})$ and benchmark set $\left\{e_{\pi}:\pi\in\mathcal{B}\right\}$. It then requests expert’s trajectory-wise annotation $a^{*,n}_{1:H}$ and updates $\mathbb{A}$ by $\ell^{n}(u)$ (Eq. (2)). At the end of round $N$, the uniform first-step mixing of $\{\pi^{n}\}_{n=1}^{N}$ is returned, which is equivalent to returning $\pi_{\hat{u}}$, where $\hat{u}:=\frac{1}{H}\sum_{n=1}^{N}u^{n}$. We provide the following performance guarantee of Algorithm 2:
Theorem 4.
If Algorithm 2 is run with a deterministic expert policy $\pi^{E}$, a policy class $\mathcal{B}$ such that realizability holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm, then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq 2R\frac{\log(B)+2\log(1/\delta)}{N}.$
Theorem 4 shows that the interactive IL Algorithm 2 matches the trajectory-wise sample complexity of behavior cloning in [4]. In contrast, prior state-of-the-art analysis of interactive IL algorithms [4, Appendix C.2] gives sample complexity results that are in general worse than behavior cloning.  [^2] For [4, Appendix C.2]’s sample complexity to improve over behavior cloning, we need $\mu H\max_{h\in[H]}\ln|\mathcal{B}_{h}|$ to be significantly smaller $R\ln|\mathcal{B}|$ (where $\mathcal{B}_{h}$ is the projection of $\mathcal{B}$ onto step $h$). This may require the strong condition that $\mu<R/H<1$ in the practically-popular parameter sharing settings ($|\mathcal{B}_{h}|=|\mathcal{B}|$).
For the proof of Theorem 4, we introduce a new notion of decoupled Hellinger estimation error:
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}:=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}$ decouples the dependency between the state sequence and the distribution of action sequence induced by the learner. Perhaps surprisingly, it is compatible with non-Markovian first-step mixture of policies, while still being well-behaved enough to be translated to a policy suboptimality guarantee, which could be of independent interest. Further details can be found in Appendix B.
3.3 Summary of Results
Table 1: Sample complexity of imitation learning algorithms to achieve $J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq\epsilon$. In the table, the cumulative cost satisfies $R\leq H$, and the recoverability factor is bounded by $1\leq\mu\leq H$.
Algorithms Annotated Trajectories Annotated States
LogLossBC [4] $O\left(\frac{R\log(B)}{\epsilon}\right)$ $O\left(\frac{RH\log(B)}{\epsilon}\right)$
LogLossDAgger [4] $O\left(\frac{\mu H\log(B)}{\epsilon}\right)$ $O\left(\frac{\mu H^{2}\log(B)}{\epsilon}\right)$
Algorithm 1 (Ours) $O\left(\frac{\mu H\log(B)}{\epsilon}\right)$
Algorithm 2 (Ours) $O\left(\frac{R\log(B)}{\epsilon}\right)$ $O\left(\frac{RH\log(B)}{\epsilon}\right)$
Table 1 summarizes the number of annotations needed to achieve $J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq\epsilon$ for both state-wise and trajectory-wise annotation settings. We include baselines from recent advancements in imitation learning, covering both offline IL (LogLossBC [4]) and interactive IL (LogLossDAgger [4]).  [4] originally does not provide sample complexity results for the state-wise annotation setting; to enable direct comparison, we convert their trajectory-wise sample complexities to state-wise sample complexities by multiplying by $H$.
For the state-wise annotation setting, DAgger with state-wise annotation oracle demonstrates significant improvements over all baselines when $\mu\ll R$. For the trajectory-wise annotation setting, DAgger with trajectory-wise annotation oracle achieves horizon-independent guarantees without the $\mu$-recoverability assumption, matching the minimax optimal performance of LogLossBC [4].
These findings provide two key clarifications: 1. Interactive IL offers potential advantage over offline IL in terms of state-wise sampling efficiency. 2. Interactive IL does not inherently suffer from suboptimal sample complexity in general compared to offline IL.
4 Related Works and Discussions
State-wise v.s. trajectory-wise annotations in IL.
The debate over using state-wise versus trajectory-wise annotations traces back to the reduction from imitation learning to supervised learning [13]. While the sample complexity of BC can be interpreted in a state-wise manner (e.g. [1, 9]), BC conventionally relies on full trajectories of expert demonstrations. Recent advancements [4] seem to settle the debate in favor of trajectory-wise annotations through more refined analysis. However, the question ”Is trajectory-wise annotation all you need?” remains unresolved in the interactive setting.
Unlike offline demonstrations, interactive annotations require querying effort. As noted by [11], “A 30-second manipulation task with 20 Hz robot control would require a human to relabel 600 state samples.” When trajectory annotation costs scale with the number of steps, annotating a single sampled state can be far more cost efficient.
Though the use of state-wise annotation appeared early in the interactive IL literature, such as in the Forward Training algorithm [13], its adoption [6] was less influential compared to DAgger [15], the interactive counterpart of BC, which popularized the convention of annotating full state sequences. Subsequent works following DAgger diverged into state-wise [9, 19, 10], trajectory-wise [20, 16, 4], and other variants [14, 1].To the best of our knowledge, we are the first to explicitly address this debate, systematically developing efficient algorithms tailored to both state-wise and trajectory-wise annotations.
First-step mixing and every-step mixing policies.
The emergence of first-step mixing originated from technical considerations. In may interactive IL methods [15, 14, 14], the returned policy was not a uniform first-step mixture but rather the best policy selected through validation. However, performing such validation in an interactive setting often requires additional expert annotations. Subsequent works [16, 9, 10, 4] circumvented the need for validation by employing a uniform first-step mixture of policies across learning rounds, thereby directly translating online regret guarantees into performance differences. While [20] implies the usage of a first-step mixture policy class, to the best of our knowledge, we are the first to establish a concrete online regret guarantee this policy class.
On the other hand, every-step mixing between the learned policy across rounds and the expert policy has been a prevalent strategy in interactive IL approaches [3, 13, 15, 14, 20]. However, when selecting the mixture policy class for each learning round, ambiguity arises regarding the choice between first-step mixing and every-step mixing [20]. For every-step mixture policies, [9] was the first to explicitly distinguish this approach from first-step mixing. In other works [16, 4], every-step mixing can be interpreted as learning $H$ separate mixture policies, one for each step within an episode.
Alternative algorithm designs and practical implementations.
Though this work follows [4] and focuses on log loss, we believe the $1/n$ rate is not exclusive to log loss. Despite requiring an additional supervision oracle, [8] suggests that trajectory-wise annotation complexity similar to Theorem 4 can be achieved using Halving [18] and 00-$1$ loss.
From an algorithmic perspective, we explored trajectory-wise annotation with first-step mixing (Algorithm 2) and state-wise annotation with every-step mixing (Algorithm 1). For trajectory-wise annotation with every-step mixing, naively learning a parameter-sharing policy may encounter a batch-summed log loss, introducing an additional $H$ factor to the sample complexity, which is undesirable. State-wise annotation with every-step mixing remains an open question for future research.
In terms of practical implementation, it is worth noting that even with oracle-efficient implementations (e.g. [9, 10]), interactive IL requires multiple computational oracle calls per learning round. In contrast, offline IL requires only a single oracle call to obtain the returned policy, which provides a clear computational advantage.
Lower bounds in interactive IL.
From an information-theoretic perspective, a line of work [17, 16, 4] provides lower bounds for imitation learning under the realizable setting and considers $\mu$-recoverability. [16] is the first to demonstrate a gap between the lower bounds of offline IL and interactive IL in trajectory-wise annotation, focusing on the non-parameter-sharing setting. [4] shifts attention back to the general setting and establishes a $\Omega(\frac{H}{\epsilon})$ sample complexity lower bound for trajectory-wise annotation.
We argue that the result in [4] also implies a $\Omega(\frac{H}{\epsilon})$ sample complexity lower bound for the state-wise annotation setting. Their proof relies on an MDP consisting only of self-absorbing states, where annotating a full trajectory is equivalent to annotating a single state. In their special MDP case ($1$-recoverable), Algorithm 1 achieves $\tilde{O}(\frac{H\log(B)}{\epsilon})$ state-wise sample complexity, which does not contradict existing lower bounds. Nonetheless, obtaining lower bounds for state-wise sample complexity for general MDPs and policy classes remains an open question.
Conclusion.
We present new results of interactive IL with log loss that matches the performance of offline IL in trajectory-wise annotation and achieves improved sample complexity in state-wise annotation. Our analysis introduced a new notion of decoupled Hellinger estimation error, a powerful tool for analyzing interactive IL algorithms. We leave the exploration of stochastic or non-realizable expert settings open for future investigation.
Appendix A Useful Distance Measures
In the following, we present 4 useful distance measures for pair of policies.
Definition 5 (Trajectory-wise $L_{1}$-divergence).
For a pair of Markovian policies $\pi$ and $\pi^{\prime}$, define their trajectory-wise $L_{1}$-divergence as
$\lambda(\pi\parallel\pi^{\prime}):=\mathbb{E}^{\pi}\mathbb{E}_{a^{\prime}_{1:H}\sim\pi^{\prime}(s_{1:H})}\left[\sum_{h=1}^{H}\mathbb{I}(a_{h}\neq a^{\prime}_{h})\right].$
$\lambda(\pi\parallel\pi^{\prime})$ is the expected total number of actions taken by $\pi^{\prime}$ that deviates from actions in trajectories induced by $\pi$. Note that $\lambda(\cdot||\cdot)$ is asymmetric, while the same concept is applied in offline and interactive IL [13, 15] with different guarantees for $\lambda(\pi\parallel\pi^{\mathrm{E}})$ and $\lambda(\pi^{\mathrm{E}}\parallel\pi)$ (Lemma 19).
Definition 6 (Trajectory-wise $L_{\infty}$-semi-metric [4]).
For a pair of Markovian policies $\pi$ and $\pi^{\prime}$, define their trajectory-wise $L_{\infty}$-semi-metric as
$\rho(\pi\parallel\pi^{\prime}):=\mathbb{E}^{\pi}\mathbb{E}_{a^{\prime}_{1:H}\sim\pi^{\prime}(s_{1:H})}\left[\mathbb{I}\left\{\exists h:a_{h}\neq a^{\prime}_{h}\right\}\right].$
$\rho(\pi\parallel\pi^{\prime})$ is the probability of any action taken by $\pi^{\prime}$ deviating from actions in trajectories induced by $\pi$, which is symmetric [4]. A bound on $\rho(\pi\parallel\pi^{\mathrm{E}})$ leads to straightforward performance difference guarantee: $J(\pi)-J(\pi^{\mathrm{E}})\leq R\cdot\rho(\pi\parallel\pi^{\mathrm{E}})$ [4] (Lemma 20).
Definition 7 (State-wise Hellinger distance).
For a pair of Markovian policies $\pi$ and $\pi^{\prime}$, define their state-wise Hellinger distance as $\mathbb{E}_{(s,h)\sim d^{\pi}}\left[D^{2}_{\text{H}}(\pi_{h}(s),\pi^{\prime}_{h}(s))\right]$.
State-wise Hellinger distance represents the expected Hellinger distance between the action distribution of $\pi$ and $\pi^{\prime}$ on $(s,h)\sim d^{\pi}$. One notable feature here is that the distance is evaluated between $\pi_{h}(\cdot\mid s)$ and $\pi^{\prime}_{h}(\cdot\mid s)$, independent of the original action $a$ taken by $\pi$ when visiting $s$. By Lemma 19, state-wise Hellinger distance can relate to trajectory-wise $L_{1}$-divergence.
Definition 8 (decoupled Hellinger distance).
For a pair of Markovian policies $\pi$ and $\pi^{\prime}$, define their decoupled Hellinger distance as $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\prime}(s_{1:H}))\right]$.
Similarly, $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\prime}(s_{1:H}))\right]$ denotes the expected Hellinger distance between the distribution of actions $\pi(s_{1:H})$ and $\pi^{\prime}(s_{1:H})$ on state sequence $s_{1:H}$ visited by $\pi$. This allows decoupled analysis for state and action sequences, which is useful for the proof of Theorem 4.
Appendix B Decoupling States and Actions: New Guarantees for DAgger
B.1 Decoupling State and Action Sequences by Decoupled Hellinger Distance
In this section, we demonstrate that similar to $D^{2}_{\text{H}}\left(\mathbb{P}^{\pi}(s_{1:H},a_{1:H}),\mathbb{P}^{\pi^{\mathrm{E}}}(s_{1:H},a_{1:H})\right)$ [4], the decoupled Hellinger distance $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right]$ that decouples states and actions is also proportionally lower bounded by a constant factor of $\rho(\pi\parallel\pi^{\mathrm{E}})$. The following two lemma shows that such relationship holds for both Markovian policies and their first-step mixtures.
Lemma 9.
Let $\pi^{\mathrm{E}}$ be a deterministic policy, and let $\pi$ be an Markovian policy. Then we have
$\frac{1}{2}\cdot\rho(\pi\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
Lemma 10.
Let $\pi^{\mathrm{E}}$ be a deterministic policy, and let $\pi_{u}$ be a first-step mixture of Markovian policies. Then we have
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
To prove these two lemmas, we first prove a special case, i.e. Lemma 11 with first-step mixture of deterministic policies. To facilitate the proofs, we introduce the following additional notations:
  • Let $\mathcal{B}^{\mathrm{Det}}$ denote the set of all deterministic, Markovian policies. We will use $\nu,\nu^{\prime}$ to denote members of $\mathcal{B}^{\mathrm{Det}}$ and $\nu_{h}(s)$ to denote the action taken by $\nu$ at $(s,h)$ when it is clear from the context.
  • Let $\mathcal{B}^{\mathrm{E}}(s_{1:h})$ represent the subset of $\mathcal{B}^{\mathrm{Det}}$ that agrees with $\pi^{\mathrm{E}}$ on the state sequence $s_{1:h}$.
  • Define $F(\nu;\nu^{\prime};\pi^{\mathrm{E}}):=\sum_{s_{1:H}}\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right],$ which evaluates the probability that $\nu^{\prime}$ disagrees with $\pi^{\mathrm{E}}$ over the distribution of $H$-step state sequences induced by $\pi$.
Lemma 11.
Let $\pi^{\mathrm{E}}$ be a deterministic Markovian policy, and let $\pi_{u}$ be a first-step mixture of deterministic Markovian policies (elements of $\mathcal{B}^{\mathrm{Det}}$). Then we have that
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
The key idea in the following proof is to lower bound $\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right],$ which reflects the asymmetric roles of the two appearances of $\pi_{u}$’s, using a symmetric formulation via function $F$ (as shown in (6)).
Proof.
Recall the first-step mixture policy in Definition 1, we start by rewriting
$\rho(\pi_{u}\parallel\pi^{\mathrm{E}})=$ $\mathbb{E}^{\pi_{u}}\left[\mathbb{I}\left\{\exists h:a_{h}\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}\right]$ (3)
$=$ $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\sum_{s_{1:H}}\mathbb{P}^{\nu}(s_{1:H},a_{1:H})\mathbb{I}\left\{\exists h:a_{h}\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}$
$=$ $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\rho(\nu\parallel\pi^{\mathrm{E}}),$
which is a weighted combination of $\rho(\nu\parallel\pi^{\mathrm{E}})$ for $\nu\in\mathcal{B}^{\mathrm{Det}}$.
Next, we turn to analyzing $D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))$. Since the deterministic expert induces a delta mass distribution over actions, we apply the elementary fact about the Hellinger distance with delta mass distribution stated in Lemma 18, yielding:
$\frac{1}{2}\parallel\pi_{u}(s_{1:H})-\pi^{\mathrm{E}}(s_{1:H})\parallel_{1}\leq D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H})).$
We recall that $\mathcal{B}^{\mathrm{E}}(s_{1:H})$ denotes the subset of $\mathcal{B}^{\mathrm{Det}}$ that agrees with $\pi^{\mathrm{E}}$ on $s_{1:H}$ and define the total weight assigned by $u$ on it as $u(\mathcal{B}^{\mathrm{E}}(s_{1:H})):=\sum_{\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:H})}u(\nu)$. Then,
$\frac{1}{2}\parallel\pi_{u}(s_{1:H})-\pi^{\mathrm{E}}(s_{1:H})\parallel_{1}=1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H})),$
which implies:
$1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H}))\leq D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H})).$ (4)
Therefore, by taking expectation over $s_{1:H}\sim\mathbb{P}^{\pi_{u}}$ in Eq. (4),
$\sum_{s_{1:H}}\mathbb{P}^{\pi_{u}}(s_{1:H})(1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H})))\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$ (5)
We now examine the expression
$\sum_{s_{1:H}}\mathbb{P}^{\pi_{u}}(s_{1:H})(1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H}))).$ (*)
Since $\pi_{u}$ is a first-step mixture of policies in $\mathcal{B}^{\mathrm{Det}}$ with weight $u$, we have $\mathbb{P}^{\pi_{u}}(s_{1:H})=\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\mathbb{P}^{\nu}(s_{1:H})$. This allows us to rewrite ()(*)( ∗ ) using the definition of $F(\nu;\nu^{\prime},\pi^{\mathrm{E}})$ as:
()(*)=( ∗ ) = $\sum_{s_{1:H}}\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\mathbb{P}^{\nu}(s_{1:H})\sum_{\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu^{\prime})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]$ (6)
$=$ $\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})\sum_{s_{1:H}}\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]$
$=$ $\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})F(\nu;\nu^{\prime};\pi^{E})$
$=$ $\frac{1}{2}\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})\left(F(\nu;\nu^{\prime};\pi^{E})+F(\nu;\nu^{\prime};\pi^{E})\right),$
where the first three equalities are by algebra and the definition of $F(\nu;\nu;\pi^{E})$. In the last equality, we use the observation that
$\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})F(\nu;\nu^{\prime};\pi^{E})=\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})F(\nu^{\prime};\nu;\pi^{E}).$
By Lemma 7 (stated below),
((*$\geq$( ∗ ) ≥ $\frac{1}{2}\cdot\frac{1}{2}\cdot\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})\left(\rho(\nu\parallel\pi^{\mathrm{E}})+\rho(\nu^{\prime}\parallel\pi^{\mathrm{E}})\right)$
$=$ $\frac{1}{2}\cdot\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\rho(\nu\parallel\pi^{\mathrm{E}})=\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}}).$
Combining the above two inequalities with Eq (5) we conclude the proof by
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq(*)\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
Lemma 12 (Symmetric Evaluation Lemma).
Given deterministic Markovian policies $\nu$, $\nu^{\prime}$, and $\pi^{\mathrm{E}}$, the following holds
$\frac{1}{2}\cdot\left(\rho(\nu\parallel\pi^{\mathrm{E}})+\rho(\nu^{\prime}\parallel\pi^{\mathrm{E}})\right)\leq F(\nu;\nu^{\prime};\pi^{\mathrm{E}})+F(\nu^{\prime};\nu;\pi^{\mathrm{E}}).$ (7)
Proof.
Recall that
$F(\nu;\nu^{\prime};\pi^{\mathrm{E}})+F(\nu^{\prime};\nu;\pi^{\mathrm{E}})=\sum_{s_{1:H}}\left(\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left[\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]\right).$
Throughout the proof, we say that $\nu$ makes a mistake at step $h$, if $\nu_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})$. Then, we can partition all state sequences $s_{1:H}\in{\mathcal{S}}^{H}$ into 4 subsets, $\mathcal{X}_{i}$, indexed by $i\in\{1,2,3,4\}$:
  1. (1)
    $\mathcal{X}_{1}:=\{s_{1:H}\;|\;\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:H})\}$;
  2. (2)
    $\mathcal{X}_{2}:=\{s_{1:H}\;|\;\exists h,s.t.\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\}$;
  3. (3)
    $\mathcal{X}_{3}:=\{s_{1:H}\;|\;\exists h,s.t.\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\}$;
  4. (4)
    $\mathcal{X}_{4}:=\{s_{1:H}\;|\;\exists h,s.t.\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\}\}$.
In words, the four subsets divide state sequences into cases where: (1) both $\nu,\nu^{\prime}$ agree with the $\pi^{\mathrm{E}}$ throughout, (2)&\&&(3) one of $\nu,\nu^{\prime}$ makes its first mistake earlier than the other, and (4) $\nu,\nu^{\prime}$ make their first mistake at the same time. It can now be easily seen that each $s_{1:H}\in{\mathcal{S}}^{H}$ lies in exactly one of such $\mathcal{X}_{i}$, and
$\mathcal{X}_{1}\cup\mathcal{X}_{2}\cup\mathcal{X}_{3}\cup\mathcal{X}_{4}={\mathcal{S}}^{H}.$
To see this, consider $h^{\text{err}}$, the first time step $h$ such that one of $\nu$ and $\nu^{\prime}$ disagree with $\pi^{\mathrm{E}}$. If $h^{\text{err}}$ does not exist, then $s_{1:H}\in\mathcal{X}_{1}$. Otherwise, $s_{1:H}$ lies in one of $\mathcal{X}_{2},\mathcal{X}_{3},\mathcal{X}_{4}$ depending on whether $\nu$ and $\nu^{\prime}$ makes mistakes at step $h^{\text{err}}$.
By definition, subset $\mathcal{X}_{1}$ denotes trajectories $s_{1:H}$ where $\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:H})$, meaning that
$\sum_{s_{1:H}\in\mathcal{X}_{1}}\left(\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left[\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]\right)=0\;.$
For the other 3 sets, i.e. $\mathcal{X}_{i}$ for $i\in\{2,3,4\}$, we can further divide each set based on the time step where the first error occurs, formally:
$\mathcal{X}_{2}^{h}:=$$\{s_{1:H}\;|\;\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\};$ (8)
$\mathcal{X}_{3}^{h}:=$$\{s_{1:H}\;|\;\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\};$
$\mathcal{X}_{4}^{h}:=$$\{s_{1:H}\;|\;\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\}.$
By definition, each pair of subsets is disjoint and $\cup_{h\in[H]}\mathcal{X}_{i}^{h}=\mathcal{X}_{i}$, for $i=2,3,4$. Note that the determination of whether $s_{1:H}\in\mathcal{X}_{i}^{h}$ only depends on $s_{1:h}$; therefore, $\mathcal{X}_{i}^{h}$ can be represented as $\tilde{\mathcal{X}}_{i}^{h}\times{\mathcal{S}}^{H-h}$, where
$\tilde{\mathcal{X}}_{i}^{h}:=\{s_{1:h}\;|\;s_{1:H}\in\mathcal{X}_{i}^{h}\}.$
Based on this observation, we have
$\sum_{s_{1:H}\in\mathcal{X}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:H})=\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h},s_{h+1:H}\in{\mathcal{S}}^{H-h}}\mathbb{P}^{\nu}(s_{1:H})=\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:h}).$
Furthermore, since deterministic policies $\nu,\nu^{\prime},\pi^{\mathrm{E}}$ agrees with each other for all $\{s_{1:h-1}|s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}\}$,
$\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:h})=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}P_{0}(s_{1})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu_{h}(s_{h}))$ (9)
$=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}P_{0}(s_{1})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu^{\prime}_{h}(s_{h}))=\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu^{\prime}}(s_{1:h}).$
This implies that
$\sum_{s_{1:H}\in\mathcal{X}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:H})=\sum_{s_{1:H}\in\mathcal{X}_{i}^{h}}\mathbb{P}^{\nu^{\prime}}(s_{1:H}),$
and therefore, summing over all $h\in[H]$,
$\sum_{s_{1:H}\in\mathcal{X}_{i}}\mathbb{P}^{\nu}(s_{1:H})=\sum_{s_{1:H}\in\mathcal{X}_{i}}\mathbb{P}^{\nu^{\prime}}(s_{1:H}).$
Now, for $\mathcal{X}_{2}$, we have
$\sum_{s_{1:H}\in\mathcal{X}_{2}}\left(\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left[\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]\right)\geq\sum_{s_{1:H}\in\mathcal{X}_{2}}\mathbb{P}^{\nu}(s_{1:H})\;,$ (10)
where we apply the fact that for all $s_{1:H}\in\mathcal{X}_{2}$, $\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})$, and dropping the second term which is nonnegative.
Similarly, for $\mathcal{X}_{3}$, we have that
$\sum_{s_{1:H}\in\mathcal{X}_{3}}\left(\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left[\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]\right)\geq\sum_{s_{1:H}\in\mathcal{X}_{3}}\mathbb{P}^{\nu^{\prime}}(s_{1:H})=\sum_{s_{1:H}\in\mathcal{X}_{3}}\mathbb{P}^{\nu}(s_{1:H}).$ (11)
Finally, for $\mathcal{X}_{4}$, we use the fact that for $s_{1:H}\in\mathcal{X}_{4}$, $\nu,\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})$ and obtain
$\sum_{s_{1:H}\in\mathcal{X}_{4}}\left(\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left[\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]\right)$ (12)
$=$ $\sum_{s_{1:H}\in\mathcal{X}_{4}}(\mathbb{P}^{\nu}(s_{1:H})+\mathbb{P}^{\nu^{\prime}}(s_{1:H}))\geq\sum_{s_{1:H}\in\mathcal{X}_{4}}\mathbb{P}^{\nu}(s_{1:H}).$
Now, we combine Eqs. (10), (11), (12) and observe that
$\sum_{s_{1:H}\in\mathcal{X}_{2}}\mathbb{P}^{\nu}(s_{1:H})+\sum_{s_{1:H}\in\mathcal{X}_{3}}\mathbb{P}^{\nu}(s_{1:H})+\sum_{s_{1:H}\in\mathcal{X}_{4}}\mathbb{P}^{\nu}(s_{1:H})\geq$ $\frac{1}{2}\sum_{s_{1:H}\in\mathcal{X}_{2}\cup\mathcal{X}_{3}\cup\mathcal{X}_{4}}\left(\mathbb{P}^{\nu}(s_{1:H})+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\right),$ (13)
which implies
$F(\nu;\nu^{\prime};\pi^{\mathrm{E}})+F(\nu^{\prime};\nu;\pi^{\mathrm{E}})\geq\frac{1}{2}\sum_{s_{1:H}\in\mathcal{X}_{2}\cup\mathcal{X}_{3}\cup\mathcal{X}_{4}}\left(\mathbb{P}^{\nu}(s_{1:H})+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\right).$ (14)
Based on the definitions of $\mathcal{X}_{2},\mathcal{X}_{3},\mathcal{X}_{4}$ and $\rho(\cdot\parallel\cdot)$,
$\sum_{s_{1:H}\in\mathcal{X}_{2}\cup\mathcal{X}_{3}\cup\mathcal{X}_{4}}\left(\mathbb{P}^{\nu}(s_{1:H})+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\right)=$ $\sum_{s_{1:H}}\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left\{\exists h:\nu_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\;\text{or}\;\nu^{\prime}_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}$ (15)
$+\sum_{s_{1:H}}\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left\{\exists h:\nu_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\;\text{or}\;\nu^{\prime}_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}$
$\geq$ $\sum_{s_{1:H}}\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left\{\exists h:\nu_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}$
$+\sum_{s_{1:H}}\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left\{\exists h:\nu^{\prime}_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}$
$=$ $\rho(\nu\parallel\pi^{\mathrm{E}})+\rho(\nu^{\prime}\parallel\pi^{\mathrm{E}}),$
where $s_{1:H}\in\mathcal{X}_{2}\cup\mathcal{X}_{3}\cup\mathcal{X}_{4}$ implies either $\nu$ or $\nu^{\prime}$ disagrees with $\pi^{\mathrm{E}}$, while the inequality relaxes the condition by splitting it into separate contributions for $\nu$ and $\nu^{\prime}$.
We conclude the proof by plugging (15) into (14).
B.1.1 Proof of Lemma 9
Proof.
The key to this proof is showing that any Markovian policy $\pi$ is equivalent, in terms of action distribution on any state sequence, to a first-step mixture of a set of deterministic Markovian policies. This leads to equivalence on trajectory distribution and decoupled Hellinger distance. To clarify further, we present the following claim, which allows us to apply guarantees for mixtures of deterministic policies in Lemma 11.
Claim 13.
For a Markovian policy $\pi$, there exists a first-step mixture of deterministic policy $\pi_{u}$ such that for any $s_{1:H}\in{\mathcal{S}}^{H}$, 1. $\pi(s_{1:H})=\pi_{u}(s_{1:H})$, and 2. $\mathbb{P}^{\pi}(s_{1:H})=\mathbb{P}^{\pi_{u}}(s_{1:H})$.
Given an MDP with finite state space size $S$ and action space size $A$, the set of all deterministic, Markovian policies, denoted by $\mathcal{B}^{\mathrm{Det}}$, contains $A^{SH}$ deterministic policies, which can be indexed by a tuple of actions $(a_{h,s})_{h\in[H],\,s\in\mathcal{S}}$.
To construct policy $\pi_{u}$, we will set the weight vector $u$ such that its weight on policy $\nu$ indexed by $(a_{h,s})_{h\in[H],s\in{\mathcal{S}}}$ as:
$u(\nu)=\prod_{h=1}^{H}\prod_{s\in{\mathcal{S}}}\pi_{h}(a_{h,s}|s)$ (16)
It can be easily verified by that $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)=1$.
We now verify the first item. By first-step mixture, we rewrite $\pi_{u}(a_{1:H}\parallel s_{1:H})$ as
$\pi_{u}(a_{1:H}\parallel s_{1:H})=$ $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\prod_{h=1}^{H}\nu(a_{h}|s_{h})$ (17)
$=$ $\sum_{(a^{\prime}_{h,s})_{h\in[H],s\in{\mathcal{S}}}}\prod_{h=1}^{H}\prod_{s\in{\mathcal{S}}}\pi_{h}(a^{\prime}_{h,s}|s)\prod_{h=1}^{H}\mathbb{I}\left[a^{\prime}_{h,s_{h}}=a_{h}\right]$
$=$ $\sum_{(a^{\prime}_{h,s})_{h\in[H],s\neq s_{h}}}\prod_{h=1}^{H}\prod_{s\neq s_{h}}\pi_{h}(a^{\prime}_{h,s}|s)\sum_{(a^{\prime}_{h,s})_{h\in[H],s=s_{h}}}\prod_{h=1}^{H}\pi_{h}(a^{\prime}_{h,s_{h}}|s_{h})\prod_{h=1}^{H}\mathbb{I}\left[a^{\prime}_{h,s_{h}}=a_{h}\right]$
$=$ $\sum_{(a^{\prime}_{h,s})_{h\in[H],s\neq s_{h}}}\prod_{h=1}^{H}\prod_{s\neq s_{h}}\pi_{h}(a^{\prime}_{h,s}|s)\prod_{h=1}^{H}\pi_{h}(a_{h}|s_{h})$
$=$ $\prod_{h=1}^{H}\pi_{h}(a_{h}|s_{h})=\pi(a_{1:H}\parallel s_{1:H}).$
Since this holds for any action sequence $a_{1:H}\in\mathcal{A}^{H}$, we derive the first part of Claim 13 that $\pi(s_{1:H})=\pi_{u}(s_{1:H})$. The second item follows from the first item in combination with Lemma 19.
We conclude that for the $\pi_{u}$ in the statement of the claim,
$\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right]=\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
Finally, the proof follows by applying Lemma 11 to $\pi_{u}$. ∎
B.1.2 Proof of Lemma 10
Proof.
By Claim 13, any Markovian policy can be viewed as a first-step mixture of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$, then any first-step mixture of Markovian policies $\pi_{u}$ can also be viewed as a first-step mixture of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$. The proof follows by applying Lemma 11. ∎
B.2 New Guarantees for DAgger with Trajectory-wise Annotation
Recall that we have defined decoupled Hellinger estimation error:
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
In the following, we first demonstrate that the performance difference between expert and the the uniform first-step mixture of any Markovian policy sequence $\{\pi^{n}\}_{n=1}^{N}$ is upper-bounded by $2R\operatorname{OnEst}^{\mathrm{Traj}}_{N}/N$, and then show the trajectory-wise sample complexity of Algorithm 2 in Theorem 4.
Lemma 14.
For any MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, and sequence of policies $\left\{\pi^{n}\right\}_{n=1}^{N}$, each of which can be Markovian or a first-step mixture of Markovian policies, their first step uniform mixture policy $\hat{\pi}$ satisfies.
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq 2R\cdot\frac{\operatorname{OnEst}^{\mathrm{Traj}}_{N}}{N}.$
Proof.
By Lemma 9 and Lemma 10, for each $\pi^{n}$, whether it is Markovian or a first-step mixture of Markovian policies, the following holds:
$\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right]\geq\frac{1}{2}\rho(\pi^{n}\parallel\pi^{\mathrm{E}}).$
Then, by the definition of $\operatorname{OnEst}^{\mathrm{Traj}}_{N}$,
$\frac{\operatorname{OnEst}^{\mathrm{Traj}}_{N}}{N}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(s^{n}_{1:H}),\pi^{\mathrm{E}}(s^{n}_{1:H}))\right]\geq\frac{1}{2N}\sum_{n=1}^{N}\rho(\pi^{n}\parallel\pi^{\mathrm{E}})=\frac{1}{2}\rho(\hat{\pi}\parallel\pi^{\mathrm{E}}),$
where we apply the fact that $\hat{\pi}$ is a first-step mixture of $\{\pi^{n}\}_{n=1}^{N}$ with uniform weights. Finally, we conclude the proof by applying Lemma 20. ∎
Theorem 15 (Theorem 4 Restated).
If Algorithm 2 is run with a deterministic expert policy $\pi^{E}$, a policy class $\mathcal{B}$ such that realizability holds, and the online learning oracle $\mathbb{A}$ set to exponential weight algorithm (see Proposition 21). Then, with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}\leq\log(B)+2\log(1/\delta),$
and furthermore, the returned $\hat{\pi}$ satisfies
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq 2R\frac{\log(B)+2\log(1/\delta)}{N}.$
Proof.
The proof closely follows Proposition C.2 in [4], tailored for another DAgger variant. However, in this case, we leverage the distribution of the state sequence $s_{1:H}$ instead of the per-step state distribution.
Observe that the log-loss functions passed through $\textbf{Alg}_{\text{Est}}$, $\ell^{n}(u)$ is of the form
$\ell^{n}(u)=\log(1/\pi_{u}(a^{n,*}_{1:H}\parallel s^{n}_{1:H}))=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a^{n,*}_{1:H}\parallel s^{n}_{1:H})}\right).$
It can be observed that $\ell^{n}$’s are 1-exp-concave. Therefore, implementing $\textbf{Alg}_{\text{Est}}$ using the exponential weights algorithm (Proposition 21) ensures that the following bound holds almost surely:
$\sum_{n=1}^{N}\log(1/\pi^{n}(a^{*,n}_{1:H}\parallel s^{n}_{1:H}))\leq\sum_{n=1}^{N}\log(1/\pi^{\mathrm{E}}(a^{*,n}_{1:H}\parallel s^{n}_{1:H}))+\log(B)=\log(B).$
Then, Lemma 22 with $x^{n}=s^{n}_{1:H}$, $y^{n}=a^{*,n}_{1:H}$, $g_{*}=\pi^{\mathrm{E}}$, and $\mathcal{H}^{n}=\{o^{n^{\prime}}\}_{n^{\prime}=1}^{n}$, where $o^{n}=\left(s_{1}^{n},a_{1}^{n},a_{1}^{*,n},\dots,s_{H}^{n},a_{H}^{n},a_{H}^{*,n}\right)$, implies that with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(s^{n}_{1:H}),\pi^{\mathrm{E}}(s^{n}_{1:H}))\right]\leq\log(B)+2\log(1/\delta).$
Finally, the second part of the theorem follows by applying Lemma 14.
Appendix C DAgger Variant with State-wise Annotation
Different from [4], instead of having access to action demonstrations at each step $h$, at round $n$, we only sample one state $(s^{n},h^{n})$ uniformly from a trajectory ($d^{\pi^{n}}$) and query for expert’s annotation.
We recall the state-wise Hellinger estimation error
$\operatorname{OnEst}^{\mathrm{State}}_{N}=\sum_{n=1}^{N}\mathbb{E}_{(s,h)\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}_{h}(s),\pi^{\mathrm{E}}_{h}(s))\right].$
In the following, we show that performance difference between $\hat{\pi}$ returned from Algorithm 1 and $\pi^{\mathrm{E}}$ can also be bounded by $\operatorname{OnEst}^{\mathrm{State}}_{N}$ in Lemma 16. Then, we show the state-wise annotation complexity of Algorithm 1 in Theorem 17.
Lemma 16.
For any MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, and sequence of policies $\left\{\pi^{n}\right\}_{n=1}^{N}$, each of which can be Markovian or a first-step mixture of Markovian policies, their first step uniform mixture policy $\hat{\pi}$ satisfies.
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq\mu H\cdot\frac{\operatorname{OnEst}^{\mathrm{State}}_{N}}{N}.$
Proof.
By Lemma 19, under the assumption of recoverability, the performance difference between $\hat{\pi}$ and the expert is bounded by
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq\mu\cdot\lambda(\hat{\pi}\parallel\pi^{\mathrm{E}}),$
where we recall the notation that
$\lambda(\pi\parallel\pi^{\mathrm{E}})=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}\mathbb{I}(a_{h}\neq\pi^{\mathrm{E}}_{h}(s_{h}))\right]=\frac{1}{2}\sum_{h=1}^{H}\mathbb{E}^{\pi}\|\pi_{h}(s_{h})-\pi^{\mathrm{E}}(s_{h})\|_{1}.$
The proof follows by upper-bounding $\sum_{n=1}^{N}\lambda(\pi^{n}\parallel\pi^{\mathrm{E}})$ by $H\cdot\operatorname{OnEst}^{\mathrm{State}}_{N}$. To this end, it suffices to show that for any Markovian policy $\pi$,
$H\cdot\mathbb{E}_{(s,h)\sim d^{\pi}}\left[D_{H}^{2}(\pi_{h}(s),\pi^{\mathrm{E}}_{h}(s))\right]\geq\frac{1}{2}\sum_{h=1}^{H}\mathbb{E}^{\pi}\|\pi_{h}(s_{h})-\pi^{\mathrm{E}}(s_{h})\|_{1}.$
Observe that $H\cdot\mathbb{E}_{(s,h)\sim d^{\pi}}\left[D_{H}^{2}(\pi_{h}(s),\pi^{\mathrm{E}}_{h}(s))\right]=\sum_{h=1}^{H}\mathbb{E}^{\pi}\left[D_{H}^{2}(\pi_{h}(s_{h}),\pi^{\mathrm{E}}_{h}(s_{h}))\right]$, we conclude the proof by applying Lemma 18 with $p=\pi_{h}(s_{h})$ and $q=\pi^{\mathrm{E}}_{h}(s_{h})$, which gives
$D^{2}_{\text{H}}(\pi_{h}(s),\pi^{\mathrm{E}}_{h}(s))\geq\frac{1}{2}\|\pi_{h}(s)-\pi^{\mathrm{E}}_{h}(s)\|_{1}.$
Theorem 17 (Theorem 3 Restated).
If Algorithm 1 is run with a deterministic expert policy $\pi^{E}$, an MDP $\mathcal{M}$ such that $(\mathcal{M},\pi^{E})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that realizability holds, and the online learning oracle $\mathbb{A}$ set to exponential weight algorithm (see Proposition 21). Then, with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{State}}_{N}\leq\log(B)+2\log(1/\delta),$
and furthermore, the returned $\hat{\pi}$ satisfies
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq\mu H\frac{\log(B)+2\log(1/\delta)}{N}.$
Proof.
Recall the each-step mixing in Definition 2, since $\bar{\pi}_{u}$ is a each-step mixing policy, where $\forall h\in[H],s\in{\mathcal{S}}$, $\bar{\pi}_{u,h}(a|s)=\sum_{\pi\in\mathcal{B}}u(\pi)\pi_{h}(a|s)$.
By using $\bar{\pi}_{u}$, the loss functions at each round that passed through $\textbf{Alg}_{\text{Est}}$, $\ell^{n}(u)$ is of the form
$\ell^{n}(u)=\log(1/\bar{\pi}_{u,h}(a^{n,*}|s^{n}))=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi_{h}(a^{n,*}|s^{n})}\right),$
which is 1-exp-concave with respect to $u$. Thus, implementing $\textbf{Alg}_{\text{Est}}$ using the exponential weights algorithm (Proposition 21) achieves:
$\sum_{n=1}^{N}\log(1/\pi^{n}_{h^{n}}(a^{*,n}\mid s^{n}))\leq\sum_{n=1}^{N}\log(1/\pi^{\mathrm{E}}_{h^{n}}(a^{*,n}\mid s^{n}))+\log(B)=\log(B).$
Then, Lemma 22 with $x^{n}=(s^{n},h^{n})$, $y^{n}=a^{*,n}$, $g_{*}=\pi^{\mathrm{E}}$, and $\mathcal{H}^{n}=\{o^{n^{\prime}}\}_{n^{\prime}=1}^{n}$, where $o^{n}=\left(s^{n},h^{n},a^{n},a^{*,n}\right)$, implies that with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{State}}_{N}=\sum_{n=1}^{N}\mathbb{E}_{(s^{n},h^{n})\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}_{h^{n}}(s^{n}),\pi^{\mathrm{E}}_{h^{n}}(s^{n}))\right]\leq\log(B)+2\log(1/\delta).$
The second part of the theorem follows by applying Lemma 16.
Appendix D Auxiliary Results
Lemma 18.
If $p,q$ are two distributions over some discrete domain $\mathcal{Z}$, and $q$ is a delta mass on an element in $\mathcal{Z}$. Then
$\frac{1}{2}\parallel p-q\parallel_{1}\leq D_{H}^{2}(p\parallel q)\leq\parallel p-q\parallel_{1}$
Lemma 19 (Performance Difference Lemma [7, 13]).
For two Markovian policies $\pi$ and $\pi^{\mathrm{E}}$ $:{\mathcal{S}}\to\Delta(\mathcal{A})$, we have
$J(\pi)-J(\pi^{\mathrm{E}})=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}A_{h}^{E}(s_{h},a_{h})\right],$
where $A_{h}^{E}(s_{h},a_{h}):=Q_{h}^{\pi^{\mathrm{E}}}(s_{h},a_{h})-V_{h}^{\pi^{\mathrm{E}}}(s_{h})$. Furthermore:
  • It holds that (recall Definition 5)
    $J(\pi)-J(\pi^{E})\leq H\cdot\lambda(\pi^{\mathrm{E}}\parallel\pi).$
  • suppose $(\operatorname{\mathcal{M}},\pi^{\mathrm{E}})$ is $\mu$-recoverable, then
    $J(\pi)-J(\pi^{E})\leq\mu\cdot\lambda(\pi\parallel\pi^{\mathrm{E}}).$
Lemma 20 (Lemma D.2. of  [4]).
For all (potentially stochastic) policies $\pi$ and $\pi^{\prime}$, it holds that
$J(\pi)-J(\pi^{\prime})\leq R\cdot\rho(\pi\parallel\pi^{\prime}).$
Proposition 21 (Proposition 3.1 of [2]).
Suppose $\{\ell^{n}(u)\}_{n=1}^{N}$ is a sequence of $\eta$-exp-concave functions from $\Delta(\mathcal{X})$ to $\mathbb{R}$. For all $x\in\mathcal{X}$, define the weights $w_{x}^{n-1}$ and probabilities $u^{n}(x)$ as follows:
$w_{x}^{n-1}=e^{-\eta\sum_{i=1}^{n-1}\ell_{i}(e_{x})},\quad u^{n}(x)=\frac{w_{x}^{n-1}}{\sum_{x^{\prime}\in\mathcal{X}}w_{x^{\prime}}^{n-1}},$
where $e_{x}$ is the $x$-th standard basis vector in $\mathbb{R}^{|\mathcal{X}|}$. Then, choosing $u^{n}=\{u^{n}(x)\}_{x\in\mathcal{X}}$ (exponential weights used with learning rate $\eta$) satisfies:
$\sum_{n=1}^{N}\ell^{n}(u^{n})\leq\min_{x\in\mathcal{X}}\sum_{n=1}^{N}\ell^{n}(e_{x})+\frac{\ln|\mathcal{X}|}{\eta}.$
Lemma 22 (Restatement of Lemma A.14 in [5]).
Under the realizbility assumption, where there exists $g_{\star}:=g_{i_{\star}}\in\mathcal{G}$ such that for all $n\in[N]$,
$y^{n}\sim g_{\star}^{n}(\cdot\mid x^{n})\mid x^{n},\mathcal{H}^{n-1},$
where $\mathcal{H}^{n-1}$ denotes all histories at the beginning of round $n$.
Then, for any estimation algorithm and any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sum_{n=1}^{N}\mathbb{E}_{n-1}\left[D_{\mathcal{H}}^{2}\left(\hat{g}^{n}(x^{n}),g_{\star}^{n}(x^{n})\right)\right]\leq\sum_{n=1}^{N}\left(\ell_{\log}^{n}(\hat{g}^{n})-\ell_{\log}^{n}(g^{n}_{\star})\right)+2\log(\delta^{-1}).$
where $\ell_{\log}^{n}\left(g\right):=\log\left(1/g\left(y^{n}\mid x^{n}\right)\right)$, and $\mathbb{E}_{n}[\cdot]:=\mathbb{E}[\cdot\mid\mathcal{H}^{n}]$.
We have the following well-known lemma for causally-conditioned probability (e.g. [21]).
Definition 23.
The causally-conditioned probability of state sequence $s_{1:H}$ given action sequence $a_{1:H-1}$, is defined as
$\mathbb{P}^{\mathcal{M}}(s_{1:H}\parallel a_{1:H-1})=P_{0}(s_{1})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})$
Lemma 24.
For any Markovian policy $\pi$,
$\mathbb{P}^{\pi}(s_{1:H},a_{1:H})=\mathbb{P}^{\mathcal{M}}(s_{1:H}\parallel a_{1:H-1})\cdot\pi(a_{1:H}\parallel s_{1:H}),$ (18)
and for any first-step mixture of Markovian policy $\pi_{u}$,
$\mathbb{P}^{\pi_{u}}(s_{1:H},a_{1:H})=\mathbb{P}^{\mathcal{M}}(s_{1:H}\parallel a_{1:H-1})\cdot{\pi_{u}}(a_{1:H}\parallel s_{1:H}).$ (19)
Proof.
Eq. (18) follows by noticing that both sides are equal to
$P_{0}(s_{1})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\pi_{h}(a_{h}\mid s_{h}).$
Eq. (19) follows by noticing that both sides are equal to
$\sum_{\nu}u(\nu)P_{0}(s_{1})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\nu_{h}(a_{h}\mid s_{h}).$