### v2

# Interactive and Hybrid Imitation Learning: Provably Beating Behavior Cloning

Abstract
Imitation learning (IL) is a paradigm for learning sequential decision-making policies from experts, leveraging offline demonstrations, interactive annotations, or both. Recent advances show that when annotation cost is tallied per trajectory, Behavior Cloning (BC)—which relies solely on offline demonstrations—cannot be improved in general, leaving limited conditions for interactive methods such as DAgger to help. We revisit this conclusion and prove that when the annotation cost is measured per state, algorithms using interactive annotations can provably outperform BC. Specifically: (1) we show that Stagger, a one‑sample‑per‑round variant of DAgger, provably beats BC under low-recovery-cost settings; (2) we initiate the study of hybrid IL where the agent learns from offline demonstrations and interactive annotations. We propose Warm-Stagger whose learning guarantee is not much worse than using either data source alone. Furthermore, motivated by compounding error and cold‑start problem in imitation learning practice, we give an MDP example in which Warm-Stagger has significant better annotation cost; (3) experiments on MuJoCo continuous‑control tasks confirm that, with modest cost ratio between interactive and offline annotations, interactive and hybrid approaches consistently outperform BC. To the best of our knowledge, our work is the first to highlight the benefit of state‑wise interactive annotation and hybrid feedback in imitation learning.
1 Introduction
Imitation learning, or learning from demonstrations, is a widely applied paradigm for learning sequential decision-making policies pomerleau1988alvinn; bojarski2016end; bansal2018chauffeurnet. In many applications, it offers a preferable alternative to reinforcement learning, as it bypasses the need for carefully designed reward functions and avoids costly exploration osa2018algorithmic; sun2017deeply.
Two prominent data collection regimes exist in imitation learning: offline and interactive. In offline imitation learning, expert demonstration data in the format of trajectories is collected ahead of time, which is a non-adaptive process that is easy to maintain. In contrast, in interactive imitation learning, the learner is allowed to query the expert for annotations in an adaptive manner ross2011reduction; ross2014reinforcement; sun2017deeply. The most basic and well-known approach for offline imitation learning is Behavior Cloning ross2010efficient; foster2024behavior, which casts the policy learning problem as a supervised learning problem that learns to predict expert actions from states. Although simple and easy to implement, offline imitation learning has the drawback that the quality of the data can be limited pomerleau1988alvinn. As a result, the trained model can well suffer from compounding error, where imperfect imitation leads the learned policy to enter unseen states, resulting in a compounding sequence of mistakes. In contrast, in interactive imitation learning, the learner maintains a learned policy over time, with the demonstrating experts providing corrective feedback on-policy, which enables targeted collection of demonstrations and improves sample efficiency.
Recent work foster2024behavior, via a sharp theoretical analysis of Behavior Cloning, shows that the sample efficiency of Behavior Cloning cannot be improved in general when measuring using the number of trajectories annotated. Interactive methods like DAgger ross2010efficient can enjoy sample complexity benefits, but so far the benefits are only exhibited in limited examples, with the most general ones in the tabular setting rajaraman2021value. This leaves open the question:
Can interaction provide sample efficiency benefit for imitation learning under a broad range of settings, especially with function approximation?
In this paper, we make progress towards this question, with a focus on the deterministically realizable setting (i.e. the expert policy $\pi^{E}$ is deterministic and is in the learner’s policy class $\mathcal{B}$). Specifically, we make the following contributions:
  1. 1.
    Motivated by the costly nature of interactive labeling on entire trajectories laskey2017comparing; mandlekar2020human, we propose to measure the cost of annotation using the number of states annotated by the demonstrating expert. We propose a general state-wise interactive imitation learning algorithm, Stagger, and show that as long as the expert can recover from mistakes at low cost in the environment ross2011reduction, it significantly improves over Behavior Cloning in terms of its number of state-wise demonstrations required.
  2. 2.
    Motivated by practical imitation learning applications where sets of offline demonstration data are readily available, we study hybrid imitation learning, where the learning agent has the additional ability to query the demonstration expert interactively to improve its performance. We design a hybrid imitation learning algorithm, Warm-Stagger, and prove that its policy optimality guarantee is not much worse than using either of the data sources alone.
  3. 3.
    Inspired by compounding error pomerleau1988alvinn and cold start problem lin2020autoregressive; oetomo2023cutting, two practical challenges in imitation learning, we provide an MDP example, for which we show hybrid imitation learning can achieve strict sample complexity savings over using either source alone, and provide simulation results that verify this theoretical claim.
  4. 4.
    We conduct experiments in MuJoCo continuous control tasks and show that if the cost of state-wise interactive demonstration is not much higher than its offline counterpart, interactive algorithms can enjoy a better cost efficiency than Behavior Cloning. Under some cost regimes and some environments, hybrid imitation learning can outperform approaches that use either source alone.
2 Preliminaries
Basic notation.
Define $n]:=\left\{1,\ldots,n\right\}$. Denote by $\Delta(\mathcal{X})$ the set of probability distributions over a set $\mathcal{X}$. For $u\in\Delta(\mathcal{X})$ and $x\in\mathcal{X}$, we denote by $u(x)$ the $x$-th coordinate of $u$ and $e_{x}$ the delta mass on $x$. We use the shorthand $x_{1:n}$ to represent the sequence $\{x_{i}\}_{i=1}^{n}$. We will frequently use the Hellinger distance to measure the difference between two distributions: $D_{H}^{2}(\mathbb{P},\mathbb{Q})=\int(\sqrt{\tfrac{d\mathbb{P}}{d\omega}}-\sqrt{\tfrac{d\mathbb{Q}}{d\omega}})^{2}d\omega$, where $\mathbb{P}$ and $\mathbb{Q}$ share a dominating measure $\omega$.
Episodic Markov decision process and agent-environment interaction.
An episodic MDP $\mathcal{M}$ is defined as a tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R},H)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:=\{P_{h}:{\mathcal{S}}\times\mathcal{A}\rightarrow\Delta({\mathcal{S}})\}_{h=1}^{H}$ denotes the transition dynamics, $\mathcal{R}:{\mathcal{S}}\times\mathcal{A}\rightarrow\Delta([0,1])$ denotes the reward distribution, and $H$ denotes episode length. Given a stationary policy $\pi:{\mathcal{S}}\rightarrow\Delta(\mathcal{A})$, we use $\pi(\cdot|s)$ to denote the action distribution of $\pi$ on $s$. A policy induces a distribution over trajectories $\tau=(s_{h},a_{h},r_{h})_{h=1}^{H}$ by first drawing the initial state $s_{1}\sim P_{0}(\varnothing)$, and then iteratively taking actions $a_{h}\sim\pi(\cdot\mid s_{h})$, receiving rewards $r_{h}\sim\mathcal{R}(s_{h},a_{h})$, and transitioning to the next state $s_{h+1}\sim P_{h}(s_{h},a_{h})$ (except at step $H$, where $P_{H}=\varnothing$). Let $\mathbb{E}^{\pi}$ and $\mathbb{P}^{\pi}$ denote expectation and probability law for $(s_{h},a_{h},r_{h})_{h=1}^{H}$ induced by $\pi$ and $\mathcal{M}$. Given $\pi$, denote by $d^{\pi}(s):=\frac{1}{H}\mathbb{P}^{\pi}(s_{h}=s)$ its state visitation distribution. The expected return of policy $\pi$ is defined as $J(\pi):=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}r_{h}\right],$ and the value functions of $\pi$ are given by $V_{h}^{\pi}(s):=\mathbb{E}^{\pi}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}\,|\,s_{h}=s\right]$, and $Q_{h}^{\pi}(s,a):=\mathbb{E}^{\pi}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}\,|\,s_{h}=s,a_{h}=a\right].$ If for policy $\pi$, step $h$, and state $s$, $\pi(\cdot\mid s)$ is the delta-mass on an action, we also sometimes slightly abuse the notation and let $\pi(s)$ denote that action.
Additional policy-related notations.
Throughout, we assume the access to an Markovian policy class $\mathcal{B}$ of finite size $B$, which contains the deterministic expert demonstrator policy $\pi^{\mathrm{E}}:{\mathcal{S}}\rightarrow\Delta(\mathcal{A})$. A $(\text{MDP},\text{Expert})$ pair $(\mathcal{M},\pi^{\mathrm{E}})$ is said to be $\mu$-recoverable if for all $h\in[H]$, $s\in{\mathcal{S}}$ and $a\in\mathcal{A}$, $Q_{h}^{\pi^{\mathrm{E}}}(s,a)-V_{h}^{\pi^{\mathrm{E}}}(s)\leq\mu$. Additionally, we assume normalized return foster2024behavior, where for any trajectory $(s_{h},a_{h},r_{h})_{h=1}^{H}$, $\sum_{h=1}^{H}r_{h}\in[0,R]$. Throughout this paper, we make the assumption that our imitation learning problem is deterministically realizable:
Assumption 1 (Deterministic Realizability).
The expert policy $\pi^{\mathrm{E}}$ is deterministic and is contained in the learner’s policy class $\mathcal{B}$.
In our algorithm and analysis, we frequently use the following “convexification” of policy class $\mathcal{B}$:
Definition 1 (Each-step Mixing of $\mathcal{B}$).
$\bar{\Pi}_{\mathcal{B}}=\{\bar{\pi}_{u,h}(a|s):=\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a|s):u\in\Delta(\mathcal{B})\}.$
An each-step mixing policy $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$ can be executed by drawing $\pi\sim u$ freshly-at-random at each step $h\in[H]$ and takes action $a\sim\pi(\cdot|s)$ (e.g. li2022efficient; li2023agnostic). Observe that $\bar{\pi}_{u}$ is a Markovian policy.
Offline imitation learning and Behavior Cloning.
In offline imitation learning, the agent is given a collection of expert trajectories $\mathcal{D}=\left\{\tau_{1},\ldots,\tau_{N_{\mathrm{off}}}\right\}$, where $\tau_{i}=(s_{i,h},a_{i,h})_{h=1}^{H}$ is the $i$-th (reward-free) trajectory, all of which are drawn iid from the trajectory distribution of expert policy $\pi^{\mathrm{E}}$. Behavior Cloning finds a policy $\pi\in\mathcal{B}$ that minimizes its log loss on expert’s actions on the seen states, i.e.,
$\hat{\pi}=\mathop{\mathrm{argmin}}_{\pi\in\mathcal{B}}\sum_{i=1}^{N_{\mathrm{off}}}\sum_{h=1}^{H}\log\frac{1}{\pi(a_{i,h}\mid s_{i,h})}.$
Recent result of foster2024behavior establishes a horizon-independent analysis of Behavior Cloning, which we recall its guarantees here:
Theorem 2 (Guarantee of BC foster2024behavior).
Suppose Assumption 1 holds, then with probability $1-\delta$, the policy returned by BC $\hat{\pi}$ satisfies:
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq\tilde{O}\left(\frac{R\log B}{N_{\mathrm{off}}}\right).$
Interactive imitation learning protocol.
In interactive IL, the learner has the ability to query the demonstration expert interactively. A first way to model interaction with expert is through a trajectory-wise demonstration oracle $\mathcal{O}^{\mathrm{Traj}}$ ross2011reduction; foster2024behavior: given a state sequence $(s_{h})_{h=1}^{H}$, return $(a_{h})_{h=1}^{H}$ such that $a_{h}=\pi^{\mathrm{E}}(s_{h})$ for all $h$. Subsequent works have considered modeling the expert as a state-wise demonstration oracle judah2014active; brantley2019disagreement; nguyen2020active; sekhari2024selective $\mathcal{O}^{\mathrm{State}}$: given a state $s_{h}$ and step $h$, return $a_{h}=\pi^{\mathrm{E}}(s_{h})$. We consider the learner interacting with the environment and demonstration oracles using the following protocol:
  • For $i=1,2,\dots$
    • Select policy $\pi^{i}$ and roll it out in $\mathcal{M}$, observing a reward-free trajectory $(s_{1},a_{1},\ldots,s_{H},a_{H})$.
    • Query the available oracle(s) to obtain expert annotations.
[…]
In practice, we expect the cost of querying $\mathcal{O}^{\mathrm{Traj}}$ to be higher than that of collecting a single offline expert trajectory laskey2017comparing. Since $H$ queries to $\mathcal{O}^{\mathrm{State}}$ can simulate one query to $\mathcal{O}^{\mathrm{Traj}}$, the cost of a single $\mathcal{O}^{\mathrm{State}}$ query should be at least $\frac{1}{H}$ the cost of $\mathcal{O}^{\mathrm{Traj}}$. Consequently, we also expect one $\mathcal{O}^{\mathrm{State}}$ query to be more expensive than obtaining an additional offline (state, expert action) pair. We denote the ratio between these two costs as $C$, where $C\geq 1$ is an application-dependent constant.  [^1] For practical settings such as human-in-the-loop learning with expert interventions mandlekar2020human; spencer2020learning, obtaining a short segment of corrective demonstrations may be cheaper than querying $\mathcal{O}^{\mathrm{State}}$ for each state therein. Here, we focus on a simplified setting and leave detailed cost modeling for such settings as interesting future work.
3 State-wise Annotation in Interactive Imitation Learning
Recent work foster2024behavior on refined analysis of Behavior Cloning (BC) casts doubt on the utility of interaction in imitation learning: when measuring sample complexity in the number of trajectories annotated, BC is minimax optimal even among interactive algorithms (foster2024behavior, Corollary 2.1 and Theorem 2.2). Although benefits of interactive approaches have been shown in specific examples, progresses so far have been sparse foster2024behavior; rajaraman2021value, with the most general results in the less-practical tabular setting rajaraman2021value. In this section, we show that interaction benefits imitation learning in a fairly general sense: when measuring sample complexity using the number of state-wise annotations, we design an interactive algorithm with sample complexity better than BC, as long as the expert policy has a low recovering cost $\mu$ in the environment.
Algorithm 1 Stagger: DAgger with State-wise annotation oracle
1:Input: MDP $\mathcal{M}$, state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$ with query budget $N_{\mathrm{int}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$.
2:for $n=1,\dots,N_{\mathrm{int}}$ do
3:  Query $\mathbb{A}$ and receive $\pi^{n}$.
4:  Execute $\pi^{n}$ and sample $s^{n}\sim d^{\pi^{n}}$. Query $\mathcal{O}^{\mathrm{State}}$ for $a^{*,n}=\pi^{\mathrm{E}}(s^{n})$.
5:  Update $\mathbb{A}$ with loss function
$\ell^{n}(\pi):=\log\left(\frac{1}{\pi(a^{*,n}|s^{n})}\right).$ (1)
6:end for
7: Output $\hat{\pi}$, a first-step uniform mixture of $\left\{\pi^{n}\right\}_{n=1}^{N_{\mathrm{int}}}$.
3.1 Interactive IL Enables Improved Sample Complexity with State-wise Annotations
Our algorithm Stagger (short for State-wise DAgger), namely Algorithm 1, interacts with the demonstration expert using a state-wise annotation oracle $\mathcal{O}^{\mathrm{State}}$. Similar to the original DAgger ross2011reduction, it requires base policy class $\mathcal{B}$ and reduces interactive imitation learning to no-regret online learning. At round $n$, it rolls out the current policy $\pi^{n}$ obtained from an online learning oracle $\mathbb{A}$ and samples $s^{n}$ from $d^{\pi^{n}}$. A classical example of $\mathbb{A}$ is the exponential weight algorithm that chooses policies from $\bar{\Pi}_{\mathcal{B}}$ (cesa2006prediction; see Proposition 38 in Appendix F). It then queries $\mathcal{O}^{\mathrm{State}}$ to get expert action $a^{*,n}$ and updates $\mathbb{A}$ with loss function $\ell^{n}(\pi)$ induced by this new example (Eq. (1)). The final policy $\hat{\pi}$ is returned as a uniform first-step mixture of the historical policies $\{\pi^{n}\}^{N_{\mathrm{int}}}_{n=1}$, i.e., sample one $\pi^{n}$ uniformly at random and execute it for the episode. In contrast to the DAgger variant analyzed in foster2024behavior, which trains a distinct policy at each step—yielding $H$ policies in total—and employs trajectory-level annotations, our algorithm utilizes parameter sharing and uses state-wise annotations.
We show the following performance guarantee of Algorithm 1 with $\mathbb{A}$ instantiated as the exponential weight algorithm:
Theorem 3.
Suppose Stagger is run with a state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$, an MDP $\mathcal{M}$ where $(\mathcal{M},\pi^{\mathrm{E}})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that deterministic realizability (Assumption 1) holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm with decision space $\Delta(\mathcal{B})$ and returns each-step mixing policies $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$. Then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq\mu H\cdot\frac{\log(B)+2\log(1/\delta)}{N_{\mathrm{int}}}.$
Theorem 3 shows that Stagger returns a policy of suboptimality $O(\frac{\mu H\log B}{N_{\mathrm{int}}})$ using $N_{\mathrm{int}}$ interactive state-wise annotations from the expert. In comparison, with the cost of $N_{\mathrm{int}}$ state-wise annotations, one can obtain $\frac{CN_{\mathrm{int}}}{H}$ trajectory-wise annotations; foster2024behavior’s analysis shows that Behavior Cloning with this number of trajectories from $\pi^{\mathrm{E}}$ returns a policy of suboptimality $O(\frac{RH\log B}{CN_{\mathrm{int}}})$; recall Theorem 2. Thus, if $C\ll\frac{R}{\mu}$, Algorithm 1 has a better cost-efficiency guarantee than Behavior Cloning.
We now sketch the proof of Theorem 3. In line with foster2024behavior, we define the online, on-policy state-wise estimation error as
$\operatorname{OnEst}^{\mathrm{State}}_{N}:=\sum_{n=1}^{N}\mathbb{E}_{s\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s),\pi^{\mathrm{E}}(\cdot\mid s))\right].$
The proof proceeds by bounding this error and translating it to the performance difference between $\hat{\pi}$ and $\pi^{\mathrm{E}}$. While our definition of estimation error is similar to (foster2024behavior, Appendix C.2), their definition requires all $H$ states per trajectory, while ours depends on the distribution over a state sampled uniformly from the rollout of policy $\pi^{n}$. This enables each labeled state to serve as immediate online feedback, fully utilizing the adaptivity of online learning. In contrast, trajectory-wise annotations may cause the online learning oracle to operate under delayed feedback joulani2013online; wan2022online, which incurs a fundamental extra factor of $H$ in state-wise sample complexity compared to our approach.
3.2 Experimental Comparison
We conduct a simple simulation study comparing the sample efficiency of log-loss Behavior Cloning foster2024behavior and Stagger in four MuJoCo todorov2012mujoco; brockman2016openai continuous control tasks with $H=1000$ and pretrained deterministic MLP experts schulman2015trust; schulman2017proximal. Considering MuJoCo’s low sensitivity to horizon length foster2024behavior, we reveal expert states one by one along consecutive trajectories for BC to allow fine-grained state-wise sample complexity comparison, while Stagger queries exactly one state per iteration by sampling from the latest policy’s rollout and updating immediately with the expert’s annotation. In Stagger, we implement the online learning oracle $\mathbb{A}$ so that it outputs a policy that approximately minimizes the log loss. In addition to log loss, we also include results with online learning oracle minimizing historical examples’ total square loss in Appendix G.2. We defer other implementation details to Appendix G.
Figure 1 shows the performance of the learned policy as a function of the number of state-wise annotations. When each interactive state-wise annotation has the same cost as an offline (state, expert action) pair ($C=1$), Stagger has superior and more stable performance than Behavior Cloning. For a given target performance (e.g., near expert-level), Stagger often requires significantly fewer state-wise annotations than BC—especially on harder tasks—though the gains are less pronounced on easier ones like Ant and Hopper. To highlight sample efficiency, we plot Stagger using only half the annotation budget of BC; despite this, it still matches or surpasses BC on several tasks, suggesting meaningful benefits from interaction when $C$ is small (e.g., $C=3$ for Walker).
Figure 1: State-wise sample complexity comparison between Behavior Cloning and Stagger. Shaded areas show the 10th–90th percentile bootstrap confidence intervals diciccio1996bootstrap over 10 runs. Stagger matches or exceeds BC with $50\%$ fewer annotations, achieving better state-wise annotation efficiency.
4 Hybrid Imitation Learning: Combining Offline Trajectory-wise and Interactive State-wise Annotations
Practical deployments of imitation learning systems often learn simultaneously from offline and interactive feedback modalities kelly2019hg; hoque2021thriftydagger: for example, in autonomous driving zhang2017query; badue2021self; zhao2023real, the learner has access to some offline expert demonstrations to start with, and also receives interactive expert demonstration feedback in trajectory segments for subsequent finetuning. Motivated by this practice, we formulate the following problem:
Hybrid Imitation Learning (HyIL): Problem Setup.
The learner has access to two complementary sources of expert supervision:
  • $N_{\mathrm{off}}$ offline expert trajectories $D_{\mathrm{off}}=\left\{(s_{i,h},a_{i,h})_{h=1}^{H},i\in[N_{\mathrm{off}}]\right\}$, sampled i.i.d. from rolling out $\pi^{E}$ in $\mathcal{M}$;
  • A state-wise annotation oracle $\mathcal{O}^{\mathrm{State}}$ that can be queried interactively up to $N_{\mathrm{int}}$ times.
[…]
We ask: can we design a HyIL algorithm with provable sample efficiency guarantee? Furthermore, can its performance surpass pure BC and pure interactive IL with the same cost budget?
Algorithm 2 Warm-Stagger: Warm-start Stagger with offline demonstrations
1:Input: MDP $\mathcal{M}$, state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$, offline expert dataset $D_{\text{off}}$ of size $N_{\mathrm{off}}$, online budget $N_{\mathrm{int}}$
2: Initialize $\mathbb{A}$ with policy class $\mathcal{B}_{\text{bc}}:=\{\pi\in\mathcal{B}:\pi(s_{h})=a_{h},\ \forall h\in[H],\forall(s_{h},a_{h})_{h=1}^{H}\in D_{\text{off}}\}$.
3:for $n=1,\dots,N_{\mathrm{int}}$ do
4:  Query $\mathbb{A}$ and receive $\pi^{n}$.
5:  Execute $\pi^{n}$ and sample $s^{n}\sim d^{\pi^{n}}$. Query $\mathcal{O}^{\mathrm{State}}$ for $a^{*,n}=\pi^{\mathrm{E}}(s^{n})$.
6:  Update $\mathbb{A}$ with loss function:
$\ell^{n}(\pi):=\log\left(\frac{1}{\pi(a^{*,n}\mid s^{n})}\right).$ (2)
7:end for
8:Output: $\hat{\pi}$, a first-step uniform mixture of $\{\pi^{1},\dots,\pi^{N}\}$.
4.1 Warm-Stagger: Algorithm and Analysis
We answer the above questions by proposing the Warm-Stagger algorithm, namely Algorithm 2. It extends Stagger to incorporate offline expert demonstrations, in that it constructs $\mathcal{B}_{\text{bc}}$, a restricted policy class that contains all policies in $\mathcal{B}$ consistent with all offline expert demonstrations (line 2). It subsequently performs online log-loss optimization on $\mathcal{B}_{\text{bc}}$ over state-action pairs collected online, where the state $s^{n}$ is obtained by rolling out $\pi^{n}$ in the MDP $\mathcal{M}$, and the action $a^{*,n}$ is annotated by the state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$. For analysis purposes, we introduce the following technical definition.
Definition 4 (Each-step policy completion).
Given a base policy class $\mathcal{B}$, define for each step $h\in[H]$
$\mathcal{B}_{h}:=\bigl\{\,\pi_{h}\;\big|\;\pi=(\pi_{1},\ldots,\pi_{H})\in\mathcal{B}\,\bigr\}.$
Then the each-step completion of $\mathcal{B}$ is defined as
$\tilde{\mathcal{B}}:=\bigl\{\,\pi=(\pi_{1},\ldots,\pi_{H})\;\big|\;\pi_{h}\in\mathcal{B}_{h}\text{ for all }h\in[H]\,\bigr\}.$
[…]
Theorem 5.
If Warm-Stagger is run with a state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$, an MDP $\mathcal{M}$ where $(\mathcal{M},\pi^{\mathrm{E}})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that deterministic realizability (Assumption 1) holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm with each-step mixing policies $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$, then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq O\left(\min\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}},\frac{\mu H\log(B_{\text{bc}}/\delta)}{N_{\mathrm{int}}}\right)\right),$
where we recall that $B\leq\tilde{B}\leq B^{H}$, and $B_{\text{bc}}:=|\mathcal{B}_{\text{bc}}|\leq B$.
Theorem 5 shows that Warm-Stagger finds a policy with suboptimality guarantee not signficantly worse than BC or Stagger: first, Behavior Cloning using the offline data only has a suboptimality of $O\left(\frac{RH\log(B/\delta)}{N_{\mathrm{off}}}\right)$ (cf. Theorem 2); second, Stagger without using offline data has a suboptimality of $O\left(\frac{\mu H\log(B/\delta)}{N_{\mathrm{int}}}\right)$ (cf. Theorem 3). We conjecture that the $\log\tilde{B}$ dependence may be sharpened to $\log B$; we leave this as an interesting open question.
Remark 6.
One may consider another baseline that naively switches between BC and Stagger based on a comparison between their bounds; however, such a baseline needs to know $R$ and $\mu$ ahead of time. In practice, we expect our Warm-Stagger to perform much better than this baseline, since it seamlessly incorporates both sources of data, and its design does not rely on theoretical bounds that may well be pessimistic.
4.2 On the Benefit of Hybrid Imitation Learning
Theorem 5 is perhaps best viewed as a fall-back guarantee for Warm-Stagger: its performance is not much worse than either of the baselines. In this section, we demonstrate that the benefit of hybrid feedback modalities can go beyond this: we construct an MDP motivated by practical challenges, in which hybrid imitation learning has a significantly better sample efficiency than both offline BC and interactive Stagger. Specifically, we prove the following theorem:
Theorem 7.
For large enough $S,H$, there exists an MDP $\mathcal{M}$ with $S$ states, and expert policy $\pi^{\mathrm{E}}$ such that:
  • With $\Omega(S)$ offline expert trajectories for BC, the learned policy is $\Omega(H)$-suboptimal;
  • With $\Omega(HS)$ interactive expert annotations for Stagger, the learned policy is $\Omega(H)$-suboptimal;
  • With $\tilde{O}\left(S/H\right)$ offline trajectories and $O(1)$ expert interactions, Warm-Stagger learns a policy $\hat{\pi}$ such that $J(\hat{\pi})=J(\pi^{E})$.
Theorem 7 suggests that when $HS\gg\max(1,C)$, Warm-Stagger achieves expert-level performance with significantly lower cost than two baselines. To see this, observe that Warm-Stagger has a total cost of $O(S+C)$, which is much smaller than $\Omega(HS)$ by BC, and $\Omega(HSC)$ by Stagger.
Figure 2: MDP construction and simulation results of algorithms with rewards assigned only in $\mathbf{E}$. We evaluate Warm-Stagger (WS) with 200, 800, 3200 offline (state, expert action) pairs. All methods are evaluated under equal total annotation cost with $C=1$. With 800 offline (state, expert action) pairs, WS significantly improves the sample efficiency over the baselines and explores $\mathbf{E^{\prime}}$ more effectively.
The MDP construction and simulation results.
We now sketch our construction of MDP $\mathcal{M}$. $\mathcal{M}$ has an episode length $H\geq 50$, $H=\Omega(\log(S))$ and action space of cardinality greater than $10H$. For each state, one of the actions is taken by the expert; the rest are “wrong” actions. We illustrate $\mathcal{M}$’s state space on the left of Figure 2; specifically, it consists of the following subsets:
  • Expert ideal states $\mathbf{E}$, where $|\mathbf{E}|=N_{0}$: this can model for example, the agent driving stably on the edge of a cliff ross2011reduction, where any incorrect action transitions the agent to the unrecoverable absorbing state set $\mathbf{B}:=\{\mathbf{b}\}$ (e.g., falling off the cliff). Taking the expert action keeps the agent in $\mathbf{E}$ with high probability $(1-\beta)$, and with a small probability $\beta$, moves the agent to $\mathbf{E^{\prime}}$ (e.g., a safe slope).
  • Unrecoverable state $\mathbf{B}=\left\{\mathbf{b}\right\}$: a special absorbing state that is unrecoverable by any action (dead).
  • Expert recoverable states $\mathbf{E^{\prime}}$: this models the agent getting off from the edge of the cliff to a safe slope. When in $\mathbf{E^{\prime}}$, taking the expert action allows the agent to return to a uniformly sampled state in $\mathbf{E}$. Taking a wrong action from $\mathbf{E^{\prime}}$ leads to $\mathbf{B^{\prime}}:=\{\mathbf{b^{\prime}}\}$ (e.g., rest area).
  • Recoverable state $\mathbf{B^{\prime}}=\left\{\mathbf{b^{\prime}}\right\}$: Not knowing how to act in $\mathbf{b^{\prime}}$ will result in the agent getting trapped in $\mathbf{b^{\prime}}$ for the episode.
We now briefly justify each algorithm’s performance as stated in Theorem 7. (1) BC only observes expert actions in $\mathbf{E}$ and $\mathbf{E^{\prime}}$, but never in $\mathbf{b^{\prime}}$. As a result, near-expert performance at test time requires high coverage over $\mathbf{E^{\prime}}$; otherwise, BC’s trained policy will likely incur compounding errors and get trapped in $\mathbf{b}^{\prime}$. (2) Stagger suffers from a cold-start problem: early policies fail to explore $\mathbf{E}$ efficiently, and incorrect actions can cause transitions into $\mathbf{b}$. Consequently, coverage over $\mathbf{E}$ grows slowly, and the policy may still fail on unannotated states in $\mathbf{E}$ even with $\Omega(HS)$ queries. (3) Warm-Stagger benefits from offline data that fully covers $\mathbf{E}$, and uses a small number of interactions to visit $\mathbf{b^{\prime}}$ and query the expert, avoiding costly exploration in $\mathbf{E^{\prime}}$ while matching expert performance.
We also conduct a simulation of the aforementioned three algorithms in a variant of the above MDP with $N_{0}=200$, $N_{1}=1000$, $H=100$, and $\beta=0.08$, using another reward function that assigns a reward of 1 only when the agent visits the states in $\mathbf{E}$. Here, we let the online learning oracle $\mathbb{A}$ optimize 0-1 loss, which is equivalent to minimizing log loss under a deterministic learner policy class and discrete actions. Figure 2 shows return and state coverage as functions of the number of expert annotations, averaged over 200 runs.
We observe that: (1) BC exhibits slow improvement, as $\mathbf{b^{\prime}}$ remains unseen (and thus unannotated) throughout training, resulting in poor performance even with substantial coverage (e.g., $80\%$) over $\mathbf{E^{\prime}}$; (2) Stagger is sample-inefficient due to slow exploration over $\mathbf{E}$ states, consistent with the cold-start intuition; (3) Warm-Stagger (WS), when initialized with limited (200) offline (state, expert action) pairs, still needs to explore $\mathbf{E}$ first before it can safely reach $\mathbf{E^{\prime}}$ without failure; and (4) Warm-Stagger with sufficient offline coverage on $\mathbf{E}$ (e.g., initialized with 3200 offline (state, expert action) pairs) directly benefits from exploring $\mathbf{b^{\prime}}$ with immediate performance gain, and enables safe and even faster exploration than the expert in $\mathbf{E^{\prime}}$.
Figure 3: Sample and cost efficiency on MuJoCo tasks. The top row shows expected return vs. number of annotations ($C=1$); the bottom row shows performance under a cost-aware setting ($C=2$). Warm-Stagger (WS) is initialized with 1/8, 1/4, or 1/2 of the total annotation budget as offline demonstrations. Specifically, WS($n$) refers to WS with offline expert trajectory demonstrations of total length $n$. For a good range of $n$’s, WS($n$) matches Stagger in sample efficiency and outperforms the baselines when $C=2$.
4.3 Hybrid IL on Continuous Control Benchmarks
Following our earlier MuJoCo-based comparison of Behavior Cloning and Stagger, we now evaluate Warm-Stagger (WS) on the same benchmarks. This experiment aims to answer: Does WS reduce total annotation cost compared to the baselines?
Based on the observation in Figure 1, we assign 400 total state-wise annotations for Hopper and Ant, and 1200 for HalfCheetah and Walker2D. For Warm-Stagger, we allocate 1/8, 1/4, or 1/2 of the total annotations to offline data, with the remainder used for interactive queries. For a fair comparison, all methods are evaluated under the same total annotation cost, with $C=1$ or $C=2$. This makes the baselines stronger, as they have full cost budget assigned to a single source.
In terms of the number of state-wise annotations ($C=1$), the results align with our theoretical findings: WS performs not significantly worse than BC or Stagger, regardless of the offline dataset size. WS still achieves performance competitive with Stagger, and even outperforms it on Ant when $C=1$. Furthermore, as shown by the purple curves, WS with appropriate offline sample size has preferable performance over 4 tasks when $C=2$, highlighting its utility in cost-aware regimes. These results confirm that Warm-Stagger reduces total annotation cost for moderate $C$.
5 Related Work
Imitation Learning with offline demonstrations
, pioneered in autonomous driving pomerleau1988alvinn, was solved by offline, state‑wise supervised learning in early works ross2010efficient; syed2010reduction and named Behavior Cloning (BC). A recent analysis by foster2024behavior employs trajectory-wise Hellinger distance to tighten the dependence of BC on the horizon at the trajectory level, although its sample complexity measured per state still grows quadratically with the horizon in the worst case. This shortcoming, often termed covariate shift or compounding error pomerleau1988alvinn, arises when imperfect imitation drives the learner to unseen states, resulting in a cascading sequence of mistakes. From a data collection perspective, this can be mitigated by noise‑injection approaches such as laskey2017dart; ke2021grasping. By leveraging additional environment interactions, generative‑adversarial IL methods ho2016generative; sun2019provably; ke2020imitation; spencer2021feedback frame learning as a two‑player game that aims to find a policy that matches expert’s state‑action visitation distributions. This setting is also known as “apprenticeship learning using inverse reinforcement learning” in earlier works  abbeel2004apprenticeship; syed2007game, which also starts from offline demonstrations and assumes the ability to interact with the environment MDP. Quantitative comparisons with these methods are beyond our scope, as they rely on extensive interactions with the MDP and access to a class of discriminator functions, while we focus on understanding the utility of state‑wise interactive annotations. This line work also include recent work of ren2024hybrid, who introduce “Hybrid Inverse Reinforcement Learning”, which leverages hybrid Reinforcement Learning to accelerate its inner loop of policy search; in contrast, our “hybrid” setting focus on utilizing heterogeneous data modalities. Recent offline imitation learning approaches chang2021mitigating; zeng2023demonstrations do not require MDP access but still require access to strong offline datasets, either with broad expert coverage or a large transition buffer. Our work assumes that interacting with the environment does not incur costs; we leave a detailed analysis that takes into account environment interaction cost as future work.
Imitation Learning with interactive demonstrations
, first proposed by ross2010efficient, allows the expert to provide corrective feedback to the learner’s action retroactively. Assuming low costs of expert recovery from mistakes, termed recoverability, DAgger ross2011reduction, and following works kim2013learning; ross2014reinforcement; sun2017deeply; cheng2018convergence; cheng2019predictor; rajaraman2021value outperform traditional BC both theoretically and empirically. However, this efficiency demands substantial annotation effort mandlekar2020human. Although DAgger ross2011reduction and some subsequent works sun2017deeply; rajaraman2021value; swamy2022minimax; foster2024behavior popularized the practice of annotating full trajectories, there has also been growing interest in state-wise annotations li2022efficient; sekhari2023selective; li2023agnostic, which appeared as early as ross2010efficient; judah2014active. In fact, practical applications of DAgger often adopt partial trajectory annotation in expert-in-the-loop mandlekar2020human; spencer2022expert; liu2022robot designs, as seen in zhang2016query; kelly2019hg; hoque2021thriftydagger; torok2025greedy, where issues such as inconsistencies caused by retroactive relabeling laskey2017comparing can be mitigated. These methods often leverage human- or machine-gated expert interventions to ensure safety during data collection zhang2017query; menda2017dropoutdagger, provide more targeted feedback menda2019ensembledagger; cui2019uncertainty, and enable learning on the fly shi2024yell. The use of selective state-wise queries aligns with our goal of promoting interactive imitation learning with efficient supervision and provable sample efficiency. We regard our contribution as providing a theoretical foundation for this increasingly popular paradigm of partial trajectory annotation.
Utilizing Offline Data for Interactive Learning.
Many practical deployments of interactive learning systems do not start from tabula rasa; instead, prior knowledge of various forms is oftentimes available. Combining offline and interactive feedback has recently gained much popularity in applications such as instruction finetuning large language models dongrlhf; ouyang2022training, and bandit machine translation nguyen2017reinforcement. Many recent theoretical works in reinforcement learning try to quantify the computational and statistical benefit of combining offline and online feedback: for example, li2023reward; tanhybrid show provable reduction of sample complexity using hybrid reinforcement learning, using novel notions of partial coverage; songhybrid shows that under some structural assumptions on the MDP, hybrid RL can bypass computational barriers in online RL kane2022computational. Many works also quantify the benefit of utilizing additional offline data sources in the contextual bandit domain; for example, oetomo2023cutting; sharma2020warm; zhang2019warm study warm-starting contextual bandits using offline bandit data and supervised learning data. While some variants of DAgger zhang2017query; hoque2021thriftydagger also operate in a hybrid setting, our work focuses on a fundamental formulation that explicitly accounts for the cost asymmetry between offline and interactive annotations settles2008active, and, to the best of our knowledge, is the first to provide a rigorous framework with provable sample efficiency guarantees.
6 Conclusion
We revisit imitation learning from the perspective of state-wise annotation. We show via the Stagger algorithm that, interaction with the demonstrating expert, with its cost properly measured, can enable provable cost efficiency gains over Behavior Cloning. We also propose Warm-Stagger that combines the benefits of offline data and interactive feedback. Our theoretical construction shows that such a hybrid method can strictly outperform both pure offline and pure interactive baselines under realistic cost models. Empirical results on the synthetic MDP support our theoretical findings, while MuJoCo experiments demonstrate the practical viability and competitive performance of our methods on continuous control tasks. Additionally, we show a trajectory-wise annotation variant of DAgger can match the sample complexity of log-loss Behavior Cloning without recoverability assumptions (Appendix E), with additional experiments (Appendix G.3).
Limitations:
Our design of imitation learning algorithm only aims at closing the gap between the performance of the expert and the trained policy; thus, the performance of our learned policy is bottlenecked by the expert’s performance. In this respect, designing imitation learners outputting policies that surpass expert performance is an important direction.
Our theory provide sample complexity guarantees for the discrete-action setting with deterministic and realizable expert. When such assumptions are relaxed, additional challenges arise simchowitz2025pitfalls. In this respect, there remains a gap between our theoretical analysis and our MuJoCo experiment results. In future work, we are interested in conducting additional experiments on discrete-action control problems (e.g., Atari) as well as language model distillation tasks.
Acknowledgments.
We thank the anonymous NeurIPS reviewers for their helpful feedback, which significantly improved the presentation of the paper. We thank National Science Foundation IIS-2440266 (CAREER) for research support.
Interactive and Hybrid Imitation Learning: Provably Beating Behavior Cloning — Supplementary Materials
Appendix A Additional Related Work
First-step mixing and each-step mixing policies.
The emergence of first-step mixing policies originated from technical considerations. In may interactive IL methods [ross2011reduction, ross2014reinforcement], the returned policy was not a uniform first-step mixture but rather the best policy selected through validation. However, performing such validation in an interactive setting often requires additional expert annotations. Subsequent works [rajaraman2021value, li2022efficient, li2023agnostic, foster2024behavior] circumvented the need for validation by employing a uniform first-step mixture of policies across learning rounds, thereby directly translating online regret guarantees into performance differences. Our Tragger algorithm (Algorithm 3 in Appendix E) also employs a first-step mixing policy at each iteration, and has sample complexity on par with behavior cloning.
On the other hand, each-step mixing between the learned policy across rounds and the expert policy has been a prevalent strategy in interactive IL approaches [daume2009search, ross2010efficient, ross2011reduction, ross2014reinforcement]. For each-step mixture policies, [li2022efficient] was the first to explicitly distinguish this approach from first-step mixing. In other works [rajaraman2021value, foster2024behavior], each-step mixing can be interpreted as learning $H$ separate mixture policies, one for each step within an episode.
Alternative algorithm designs and practical implementations.
Though this work follows [foster2024behavior] and focuses on log loss, we believe the $1/n$-rate is not exclusive to log loss. Despite requiring an additional supervision oracle, [le2018hierarchical] suggests that trajectory-wise annotation complexity similar to Theorem 3 (and Theorem 27) can be achieved using Halving [shalev2011online] and 0-11 loss.
From an algorithmic perspective, we explored trajectory-wise annotation with first-step mixing (Algorithm 3 in Appendix E) and state-wise annotation with each-step mixing (Algorithm 1). For trajectory-wise annotation with each-step mixing, naively learning a parameter-sharing policy may encounter a batch-summed log loss, introducing an (undesirable) additional $H$ factor to the sample complexity. Analyzing state-wise annotation with each-step mixing remains an open question for future research.
In terms of practical implementation, it is worth noting that even with oracle-efficient implementations (e.g. [li2022efficient, li2023agnostic]), interactive IL may require multiple supervised learning oracle calls per iteration. In contrast, offline IL requires only a single oracle call to obtain the returned policy, which provides a clear computational advantage. We also note that real-world experts can be suboptimal; in some applications it may be preferable to combine imitation and reinforcement learning signals (e.g., [ross2014reinforcement, sun2018truncated, amortila2022few]).
Lower bounds in interactive imitation learning.
From an information-theoretic perspective, a line of work [rajaraman2020toward, rajaraman2021value, foster2024behavior] provides lower bounds for imitation learning under the realizable setting and considers $\mu$-recoverability. [rajaraman2021value] is the first to demonstrate a gap between the lower bounds of offline IL and interactive IL in trajectory-wise annotation, focusing on the tabular and non-parameter-sharing setting. [foster2024behavior] establishes a $\Omega\left(\frac{H}{\epsilon}\right)$ sample-complexity lower bound for trajectory-wise annotation in the parameter-sharing setting.
We observe that the proof of  [foster2024behavior, Theorem 2.2] also implicitly implies a $\Omega(\frac{H}{\epsilon})$ sample complexity lower bound for the state-wise annotation setting. Their proof relies on an MDP consisting only of self-absorbing states, where annotating a full trajectory gives the same amount of information as annotating a single state. In that MDP (which is $1$-recoverable), Algorithm 1 achieves $\tilde{O}(\frac{H\log(B)}{\epsilon})$ state-wise sample complexity, which does not contradict this lower bound. Nonetheless, obtaining lower bounds for state-wise sample complexity for general MDPs, policy classes, and general recoverability constants remains an open question.
Appendix B Proof for Stagger
We first present 2 useful distance measures for pair of policies.
Definition 8 (Trajectory-wise $L_{1}$-divergence).
For a pair of Markovian policies $\pi$ and $\pi^{\prime}$, define their trajectory-wise $L_{1}$-divergence as
$\lambda(\pi\parallel\pi^{\prime}):=\mathbb{E}^{\pi}\mathbb{E}_{a^{\prime}_{1:H}\sim\pi^{\prime}(\cdot|s_{1:H})}\left[\sum_{h=1}^{H}\mathbb{I}(a_{h}\neq a^{\prime}_{h})\right].$
$\lambda(\pi\parallel\pi^{\prime})$ is the expected total number of actions taken by $\pi^{\prime}$ that deviates from actions in trajectories induced by $\pi$. Note that $\lambda(\cdot||\cdot)$ is asymmetric, while the same concept is applied in offline and interactive IL [ross2010efficient, ross2011reduction] with different guarantees for $\lambda(\pi\parallel\pi^{\mathrm{E}})$ and $\lambda(\pi^{\mathrm{E}}\parallel\pi)$ (Lemma 36).
Definition 9 (State-wise Hellinger distance).
For a pair of policies $\pi$ and $\pi^{\prime}$, define their state-wise Hellinger distance as $\mathbb{E}_{s\sim d^{\pi}}\left[D^{2}_{\text{H}}(\pi(\cdot\mid s),\pi^{\prime}(\cdot\mid s))\right]$.
State-wise Hellinger distance represents the expected Hellinger distance between the action distribution of $\pi$ and $\pi^{\prime}$ on $s\sim d^{\pi}$. One notable feature here is that the distance is evaluated between $\pi(\cdot\mid s)$ and $\pi^{\prime}(\cdot\mid s)$, independent of the original action $a$ taken by $\pi$ when visiting $s$. By Lemma 36, state-wise Hellinger distance can relate to trajectory-wise $L_{1}$-divergence.
In the following, we show that the performance difference between the policy $\hat{\pi}$ returned by Stagger (Algorithm 1) and the expert policy $\pi^{\mathrm{E}}$ can be bounded by the state-wise Hellinger estimation error:
$\operatorname{OnEst}^{\mathrm{State}}_{N}:=\sum_{n=1}^{N}\mathbb{E}_{s\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s),\pi^{\mathrm{E}}(\cdot\mid s))\right],$
where $\pi^{n}(\cdot\mid s)$ and $\pi^{E}_{h}(\cdot\mid s)$ denote the action distributions over $\mathcal{A}$ produced by the policies $\pi^{n}$ and $\pi^{E}$ at state $s$.
We first prove this in Lemma 10, and then prove the state-wise annotation complexity of Algorithm 1 in Theorem 11.
Different from [foster2024behavior], where access to full action demonstrations is assumed, we consider a more restrictive model where, at each round $n$, only a single state $(s^{n},h^{n})$ from the trajectory induced by $\pi^{n}$ is sampled and annotated by the expert.
Lemma 10.
For any MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, and sequence of policies $\left\{\pi^{n}\right\}_{n=1}^{N}$, then $\hat{\pi}$, the first-step uniform mixture of $\{\pi^{1},\dots,\pi^{N}\}$ satisfies:
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq\mu H\cdot\frac{\operatorname{OnEst}^{\mathrm{State}}_{N}}{N}.$
Proof.
By Lemma 36, under the assumption of recoverability, the performance difference between $\hat{\pi}$ and the expert is bounded by
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq\mu\cdot\lambda(\hat{\pi}\parallel\pi^{\mathrm{E}}),$
where we recall the notation that
$\lambda(\pi\parallel\pi^{\mathrm{E}})=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}\mathbb{I}(a_{h}\neq\pi^{\mathrm{E}}(s_{h}))\right]=\frac{1}{2}\sum_{h=1}^{H}\mathbb{E}^{\pi}\|\pi(s_{h})-\pi^{\mathrm{E}}(s_{h})\|_{1}.$
The proof follows by upper-bounding $\sum_{n=1}^{N}\lambda(\pi^{n}\parallel\pi^{\mathrm{E}})$ by $H\cdot\operatorname{OnEst}^{\mathrm{State}}_{N}$. To this end, it suffices to show that for any stationary policy $\pi$,
$H\cdot\mathbb{E}_{s\sim d^{\pi}}\left[D_{H}^{2}(\pi(\cdot\mid s),\pi^{\mathrm{E}}(\cdot\mid s))\right]\geq\frac{1}{2}\sum_{h=1}^{H}\mathbb{E}^{\pi}\|\pi(\cdot\mid s)-\pi^{\mathrm{E}}(\cdot\mid s)\|_{1}.$
Observe that $H\cdot\mathbb{E}_{s\sim d^{\pi}}\left[D_{H}^{2}(\pi(\cdot\mid s),\pi^{\mathrm{E}}(\cdot\mid s))\right]=\sum_{h=1}^{H}\mathbb{E}^{\pi}\left[D_{H}^{2}(\pi(\cdot\mid s_{h}),\pi^{\mathrm{E}}(\cdot\mid s_{h}))\right]$, we conclude the proof by applying Lemma 35 with $p=\pi(s_{h})$ and $q=\pi^{\mathrm{E}}(s_{h})$, which gives
$D^{2}_{\text{H}}(\pi(\cdot\mid s_{h}),\pi^{\mathrm{E}}(\cdot\mid s_{h}))\geq\frac{1}{2}\|\pi(\cdot\mid s_{h})-\pi^{\mathrm{E}}(\cdot\mid s_{h})\|_{1}.$
Theorem 11 (Theorem 3 Restated).
If Stagger (Algorithm 1) is run with a state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$, an MDP $\mathcal{M}$ where $(\mathcal{M},\pi^{\mathrm{E}})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that deterministic realizability (Assumption 1) holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm, then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{State}}_{N_{\mathrm{int}}}\leq\log(B)+2\log(1/\delta),$
and furthermore, the returned $\hat{\pi}$ satisfies
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq\mu H\frac{\log(B)+2\log(1/\delta)}{N_{\mathrm{int}}}.$
Proof.
Recall the each-step mixing in Definition 1, since $\bar{\pi}_{u}$ is a each-step mixing policy, where $\forall h\in[H],s\in{\mathcal{S}}$, $\bar{\pi}_{u}(a|s)=\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a|s)$.
By using $\bar{\pi}_{u}$, the loss functions at each round that passed through online learning oracle $\mathbb{A}$, $\ell^{n}(\pi)$ is of the form
$\ell^{n}(\pi)=\log(1/\bar{\pi}_{u}(a^{n,*}|s^{n}))=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a^{n,*}|s^{n})}\right),$
which is 1-exp-concave with respect to $u$. Thus, implementing $\mathbb{A}$ using the exponential weights algorithm (Proposition 38) achieves:
$\sum_{n=1}^{N_{\mathrm{int}}}\log(1/\pi^{n}_{h^{n}}(a^{*,n}\mid s^{n}))\leq\sum_{n=1}^{N_{\mathrm{int}}}\log(1/\pi^{\mathrm{E}}_{h^{n}}(a^{*,n}\mid s^{n}))+\log(B)=\log(B).$
Then, Lemma 39, a standard online-to-batch conversion argument with $x^{n}=(s^{n},h^{n})$, $y^{n}=a^{*,n}$, $g_{*}=\pi^{\mathrm{E}}$, and $\mathcal{H}^{n}=\{o^{n^{\prime}}\}_{n^{\prime}=1}^{n}$, where $o^{n}=\left(s^{n},h^{n},a^{n},a^{*,n}\right)$, implies that with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{State}}_{N_{\mathrm{int}}}=\sum_{n=1}^{{N_{\mathrm{int}}}}\mathbb{E}_{(s^{n},h^{n})\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}_{h^{n}}(\cdot\mid s^{n}),\pi^{\mathrm{E}}_{h^{n}}(\cdot\mid s^{n}))\right]\leq\log(B)+2\log(1/\delta).$
The second part of the theorem follows by applying Lemma 10. ∎
Appendix C Proof for Warm-Stagger
In this section, we analyze the guarantees of Warm-Stagger under the realizable and deterministic expert assumption. We show that all intermediate policies, as well as the final returned mixture policy, enjoy small Hellinger distance to the expert’s trajectory distribution, due to their agreement on the offline dataset. Our analysis builds on maximum likelihood estimator (MLE) generalization guarantees under log-loss minimization, and leverages the fact that each each-step mixing policy in Warm-Stagger can be viewed as a first-step policy mixture. The following claim shows that any Markovian policy $\pi$ can be represented as a first-step mixing over deterministic Markovian policies from a finite class.
Lemma 12.
Let $\bar{\pi}_{u}$ be a each-step mixing policy that, at each step $h\in[H]$, samples a base policy $\pi\in\mathcal{B}$ independently according to $u\in\Delta(\mathcal{B})$. Then, its induced trajectory distribution is equivalent to a first-step mixing the each-step policy completion of $\mathcal{B}$, denoted by $\tilde{\mathcal{B}}:=\left\{(\pi_{1},\dots,\pi):\pi\in\mathcal{B}\right\}$ (see Definition 4).
Proof.
Let $\nu=(\pi_{1},\dots,\pi_{H})\in\tilde{\mathcal{B}}$. Define $u(\nu):=\prod_{h=1}^{H}u(\pi_{h})$, which is a valid distribution over $\tilde{\mathcal{B}}$. Consider the joint action distribution under $\bar{\pi}_{u}$, which samples $\pi_{h}\sim u$ independently for each step and executes $a_{h}\sim\pi_{h}(\cdot\mid s_{h})$. The resulting conditional distribution over actions given the state sequence is
$\bar{\pi}_{u}(a_{1:H}\mid s_{1:H})=\prod_{h=1}^{H}\left(\sum_{\pi_{h}\in\mathcal{B}}u(\pi_{h})\,\pi_{h}(a_{h}\mid s_{h})\right).$
Under the first-step mixture policy $\pi_{u(\nu)}$ over $\tilde{\mathcal{B}}$, a full tuple $\nu=(\pi_{1},\dots,\pi_{H})$ is sampled once from $u(\nu)$, and actions are drawn as $a_{h}\sim\pi_{h}(\cdot\mid s_{h})$. The resulting action distribution is
$\pi_{u(\nu)}(a_{1:H}\mid s_{1:H})=\sum_{\nu\in\tilde{\mathcal{B}}}u(\nu)\prod_{h=1}^{H}\pi_{h}(a_{h}\mid s_{h}).$
Expanding the sum yields
$\sum_{(\pi_{1},\dots,\pi_{H})}\left(\prod_{h=1}^{H}u(\pi_{h})\pi_{h}(a_{h}\mid s_{h})\right)=\prod_{h=1}^{H}\sum_{\pi_{h}\in\mathcal{B}}u(\pi_{h})\,\pi_{h}(a_{h}\mid s_{h}),$
by the distributive property and independence of the product.
[…]
Lemma 13.
Let $\mathcal{B}_{\text{bc}}:=\{\pi\in\mathcal{B}:\pi(s)=\pi^{E}(s)\ \forall s\in D_{\text{off}}\}$ be the set of policies that agree with the expert on the offline dataset of $N_{\mathrm{off}}$ expert trajectories. Assume the expert $\pi^{E}$ is deterministic and realizable. Then, with probability at least $1-\delta$, for all $\pi^{i}$ generated in Warm-Stagger (Algorithm 2), it holds that:
$D_{H}^{2}(P_{\pi^{i}},P_{\pi^{E}})\leq O\left(\frac{\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right).$
Furthermore, the returned policy $\hat{\pi}$ satisfies:
$D_{H}^{2}(P_{\hat{\pi}},P_{\pi^{E}})\leq O\left(\frac{H\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right).$
Proof.
Let $\mathcal{B}_{\text{bc}}$ denote the realizable class of policies that agree with $\pi^{E}$ on the offline dataset. By the MLE generalization bound [foster2024behavior], for any single policy $\pi\in\mathcal{B}_{\text{bc}}$, we have with probability at least $1-\delta$:
$D_{H}^{2}(P_{\pi},P_{\pi^{E}})\leq O\left(\frac{\log(B/\delta)}{N_{\mathrm{off}}}\right).$
However, in Warm-Stagger, policies $\pi^{i}$ are not selected directly from $\mathcal{B}_{\text{bc}}$, but are instead mixtures over $\mathcal{B}_{\text{bc}}$ at each time step. A full trajectory is therefore sampled by first choosing a policy $\pi^{h}$ at each step $h\in[H]$, inducing an effective policy over sequences from $\tilde{\mathcal{B}}_{\text{bc}}$. Therefore, by appling $\tilde{\mathcal{B}}$ for Theorem 2, the BC guarantee holds for all $\pi\in\tilde{\mathcal{B}}$, with a factor of $\log|\tilde{\mathcal{B}}|\leq H\log B$ in the worst case instead of $\log B$. Therefore, by Lemma 12, each $\pi^{i}$ can be viewed as first-step mixing and satisfies:
$D_{H}^{2}(P_{\pi^{i}},P_{\pi^{E}})\leq O\left(\frac{H\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right).$
This remains a worst-case bound. If the per-step base policies are drawn from a factored structure (e.g., $\pi=(\pi_{1},\dots,\pi_{H})$ with $\pi_{h}\in\mathcal{B}_{\text{bc}}$), and the support is shared across steps, the effective covering number can be much smaller, reducing the overhead back to $\log B$.
Finally, the returned policy $\hat{\pi}$ is a first-step mixing of $\{\pi_{1},\dots,\pi_{N}\}$, and thus the same Hellinger bound carries over by convexity. ∎
Theorem 14 (Theorem 5 Restated).
If Algorithm 2 is run with a deterministic expert policy $\pi^{E}$, an MDP $\mathcal{M}$ such that $(\mathcal{M},\pi^{E})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that realizability holds, and the online learning oracle $\mathcal{O}$ set as the exponential weight algorithm, then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq O\left(\min\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}},\frac{\mu H\log(B_{\text{bc}}/\delta)}{N_{on}}\right)\right),$
Proof.
By Lemma 13, with probability at least $1-\delta/2$, for all $\pi\in\{\pi_{1},\cdots,\pi_{N}\}$, it satisfies that $D_{H}^{2}(P_{\pi},P_{\pi^{\mathrm{E}}})\leq O\left(\frac{\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right)$. This implies the first-step uniform mixture of $\{\pi_{1},\cdots,\pi_{N}\}$ satisfies
$D_{H}^{2}(P_{\hat{\pi}},P_{\pi^{\mathrm{E}}})=\frac{1}{N}\sum_{i=1}^{N}D_{H}^{2}(P_{\pi^{i}},P_{\pi^{\mathrm{E}}})\leq O\left(\frac{\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right)$
By invoking Theorem [foster2024behavior][2.1], we have,
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq O\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right)$
For the second half of the proof, we notice that by definition $\pi^{\mathrm{E}}\in\mathcal{B}_{\text{bc}}$. Then, by applying Theorem 3, with probability at least $1-\delta/2$, the returned $\hat{\pi}$ satisfies
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq O\left(\frac{\mu H\log(B_{\text{bc}}/\delta)}{N_{\mathrm{int}}}\right).$
Together, we conclude our proof by applying union bound. ∎
Appendix D Proof for Theorem 7
We formally define the MDP $\mathcal{M}$ from Section 4.2, where the expert policy $\pi^{\mathrm{E}}$ is deterministic and the transition dynamics are time-homogeneous across all steps $h\in[H]$.
  • State Space ${\mathcal{S}}=\mathbf{E}\cup\mathbf{E^{\prime}}\cup\mathbf{B}\cup\mathbf{B^{\prime}}$, where:
    • $\mathbf{E}$: ideal expert states, $|\mathbf{E}|=N_{0}$;
    • $\mathbf{E^{\prime}}$: recoverable expert states, $|\mathbf{E^{\prime}}|=N_{1}$;
    • $\mathbf{B}=\{\mathbf{b}\}$: absorbing failure state (unrecoverable);
    • $\mathbf{B^{\prime}}=\{\mathbf{b^{\prime}}\}$: recoverable reset state.
  • Action Space $\mathcal{A}$: $A=|\mathcal{A}|$ discrete actions. For each state $s\in{\mathcal{S}}$, there is a unique action $\pi^{\mathrm{E}}(s)$ taken by the expert.
  • Episode length $H$.
  • Initial State Distribution $\rho$:
    $\rho(s)=\frac{1}{(1+\beta)N_{0}}\;\text{for all}\;s\in\mathbf{E},\quad\rho(s)=\frac{\beta}{(1+\beta)N_{1}}\;\text{for all}\;s\in\mathbf{E}^{\prime}.$
  • Transition Dynamics:
    • $s\in\mathbf{E}$:
        $a=\pi^{\mathrm{E}}(s)$: with probability $1-\beta$, transitions to a uniformly random $s^{\prime}\in\mathbf{E}$; with probability $\beta$, transitions to a uniformly random $s^{\prime}\in\mathbf{E^{\prime}}$ ;
        $a\neq\pi^{\mathrm{E}}(s)$: transitions to $\mathbf{b}$.
    • $s\in\mathbf{E^{\prime}}$:
        $a=\pi^{\mathrm{E}}(s)$: transitions to a uniformly random $s^{\prime}\in\mathbf{E}$;
        $a\neq\pi^{\mathrm{E}}(s)$: transitions to $\mathbf{b^{\prime}}$.
    • $s\in\mathbf{B}=\left\{\mathbf{b}\right\}$: absorbing for all actions. Specifically, $P(s^{\prime}=\mathbf{b}|s=\mathbf{b},a)=1$.
    • $s\in\mathbf{B^{\prime}}=\left\{\mathbf{b}^{\prime}\right\}$:
        $a=\pi^{\mathrm{E}}(\mathbf{b^{\prime}})$: transitions to a uniformly random $s^{\prime}\in\mathbf{E}$;
        $a\neq\pi^{\mathrm{E}}(\mathbf{b^{\prime}})$: remains in $\mathbf{b^{\prime}}$.
  • Reward Function: For theoretical analysis, we consider the following reward function $R_{1}$:
    $R_{1}(s,a)=\begin{cases}1&\text{if }s\in\mathbf{E}\cup\mathbf{E^{\prime}}\\ 1&\text{if }s=\mathbf{b^{\prime}}\text{ and }a=\pi^{\mathrm{E}}(s)\\ 0&\text{otherwise}\end{cases}$
  • Specification of Parameters: In the following proofs, we consider $H\geq 50$, $H\geq\frac{5}{4}\log(10N_{0})$, $A\geq 10H$, $\beta=\frac{8}{H-8}$, $N_{1}\geq 500$, and $N_{1}\geq 160N_{0}$.
We also make the following assumption on the output policy of our learning algorithms (BC, Stagger, and Warm-Stagger). At any stage of learning, denote the set of states that are annotated by the expert by ${\mathcal{S}}_{\text{annotated}}$. Given the annotated (state, expert action) pairs, the learner calls some offline or online learning oracle $\mathbb{A}$ to return a policy $\pi$. We require $\pi$’s behavior as follows:
$\pi(\cdot\mid s)=\begin{cases}\pi^{\mathrm{E}}(\cdot\mid s),&s\in{\mathcal{S}}_{\text{annotated}},\\ (\frac{1}{A},\ldots,\frac{1}{A}),&s\notin{\mathcal{S}}_{\text{annotated}}.\end{cases}$
In other words, $\pi$ follows the expert’s action whenever such information is available; otherwise, it takes an action uniformly at random.
We next make a simple observation that by the construction of $\mathcal{M}$, the expert policy’s state-visitation distribution in $\mathcal{M}$ is stationary over all time steps:
Observation 1.
For MDP $\mathcal{M}$, the expert policy $\pi^{E}$’s visitation distribution at time step $h$, $d_{h}^{\pi^{E}}$ equals $\rho$, for all $h\in[H]$.
Proof.
The initial distribution gives:
$\rho(s)=\begin{cases}\frac{1}{N_{0}(1+\beta)},\;&s\in\mathbf{E},\\ \frac{\beta}{N_{1}(1+\beta)},\;&s\in\mathbf{E^{\prime}},\\ 0,&\text{otherwise}.\end{cases}$
Under $\pi^{\mathrm{E}}$, the induced transition kernel on $\mathbf{E}\cup\mathbf{E^{\prime}}$ satisfies:
$P(s^{\prime}|s,\pi^{E}(s))=\begin{cases}\frac{1-\beta}{N_{0}},&s\in\mathbf{E},\;s^{\prime}\in\mathbf{E},\\ \frac{\beta}{N_{1}},&s\in\mathbf{E},\;s^{\prime}\in\mathbf{E^{\prime}},\\ \frac{1}{N_{0}},&s\in\mathbf{E^{\prime}},\;s^{\prime}\in\mathbf{E},\\ 0,&\text{otherwise}.\end{cases}$
For any fixed $s\in\mathbf{E}$, using the kernel above,
$\sum_{s^{\prime}}\rho(s^{\prime})P(s\mid s^{\prime},\pi^{E}(s^{\prime}))=\sum_{s^{\prime}\in\mathbf{E}}\frac{1}{N_{0}(1+\beta)}\cdot\frac{1-\beta}{N_{0}}+\sum_{s^{\prime}\in\mathbf{E^{\prime}}}\frac{\beta}{N_{1}(1+\beta)}\cdot\frac{1}{N_{0}}=\frac{1}{N_{0}(1+\beta)}=\rho(s).$
Similarly, for any fixed $s\in\mathbf{E^{\prime}}$,
$\sum_{s^{\prime}}\rho(s^{\prime})P(s\mid s^{\prime},\pi^{E}(s^{\prime}))=\sum_{s^{\prime}\in\mathbf{E}}\frac{1}{N_{0}(1+\beta)}\cdot\frac{\beta}{N_{1}}=\frac{\beta}{N_{1}(1+\beta)}=\rho(s).$
Thus, for any $s\in\mathbf{E}\cup\mathbf{E^{\prime}}$,
$\rho(s)=\sum_{s^{\prime}}\rho(s^{\prime})P(s\mid s^{\prime},\pi^{E}(s^{\prime})).$
Hence $\rho$ is a stationary distribution for the Markov chain induced by $\pi^{E}$ (and $\mathbf{B},\mathbf{B^{\prime}}$ are never reached under $\pi^{E}$). Since $d^{\pi^{E}}_{1}=\rho$ and the dynamics are time-homogeneous, induction gives $\forall h\in[H]$, $d^{\pi^{E}}_{h}=\rho$, i.e.,
$d^{\pi^{E}}_{h}(s)=\begin{cases}\frac{1}{N_{0}(1+\beta)},\;&s\in\mathbf{E},\\ \frac{\beta}{N_{1}(1+\beta)},\;&s\in\mathbf{E^{\prime}},\\ 0,&\text{otherwise}.\end{cases}$
Theorem 15 (Restatement of Theorem 7).
To achieve smaller than $\frac{H}{2}$ suboptimality compared to expert in MDP $\mathcal{M}$ with probability $\frac{1}{2}$:
  • Behavior Cloning (BC) using offline expert trajectories requires
    $N_{\mathrm{off}}=\Omega(N_{1})\quad\text{with total annotation cost }\Omega(HN_{1}).$
  • Stagger that collects interactive state-wise annotations requires
    $N_{\mathrm{int}}=\Omega(HN_{0})\quad\text{with total annotation cost }\Omega(CHN_{0}).$
In contrast, Warm-Stagger learns a policy that achieves expert performance with probability at least $\frac{1}{2}$, using
$N_{\mathrm{off}}$ $=O(\frac{N_{0}}{H}\log(N_{0}))\quad\text{expert trajectories, and}$ (3)
$N_{\mathrm{int}}$ $\leq 3\quad\text{interactive annotations,}$
$\text{with total annotation cost }\tilde{O}(N_{0}+C).$
Proof.
The proof is divided into three parts:
First, by Lemma 16, we show that in $\mathcal{M}$, Behavior Cloning requires $\Omega(HN_{1})$ expert trajectories to achieve suboptimality $H/2$ with probability $\frac{1}{2}$.
Next, in Lemma 20, we show that Stagger, which rolls out the learner policy and queries the expert on only one state sampled uniformly from its learned policy’s rollout, requires $N_{\mathrm{int}}=\Omega(HN_{0})$ interactive annotations to achieve suboptimality no greater than $H/2$ with probability $\frac{1}{2}$.
Finally, by Lemma 22, we demonstrate that Warm-Stagger achieves expert performance using $O(\frac{N_{0}}{H}\log(N_{0}))$ offline expert tarjectories and 3 interactive annotations with probability $\frac{1}{2}$. ∎
D.1 Lower Bound for Behavior Cloning
Throughout this subsection, we denote by $\mathbf{E}_{\text{annotated}}^{\prime}$ the set of states in $\mathbf{E^{\prime}}$ that are visited and annotated by the expert’s $N_{\mathrm{off}}$ offline trajectories.
Lemma 16 (BC suboptimality lower bound).
Consider the MDP $\mathcal{M}$ and the expert policy $\pi^{\mathrm{E}}$ constructed as above. If Behavior Cloning uses
$N_{\mathrm{off}}<\frac{N_{1}}{160}$
expert trajectories, then with probability at least $\frac{1}{2}$, the suboptimality of its returned policy $\hat{\pi}$ is lower bounded by:
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\geq\frac{H}{2}.$
Proof.
[…]
In the following, we show that an insufficient number of expert trajectories leads to small $\mathbf{E}_{\text{annotated}}^{\prime}$ and poor coverage on states in $\mathbf{E^{\prime}}$. This, in turn, causes the BC-learned policy to frequently fail to recover and get trapped in the absorbing bad state $\mathbf{b}^{\prime}$, incurring a large suboptimality compared to the expert.
By Lemma 17, when $N_{\mathrm{off}}$ is below the stated threshold, at most $1/10$ of the states in $\mathbf{E^{\prime}}$ is annotated by expert trajectories with probability $\frac{1}{2}$. Suppose we roll out the policy $\tilde{\pi}$ in $\mathcal{M}$; let $\tau$ be the first step such that $s_{\tau}\in\mathbf{E^{\prime}}$.
By Lemma 18,
$\Pr_{\mathcal{M},\tilde{\pi}}(\tau\leq H/5)\geq 0.79,\;\text{and}\;\Pr_{\mathcal{M},\tilde{\pi}}\big(s_{\tau}\notin\mathbf{E}_{\text{annotated}}^{\prime}\mid\tau\leq H/5\big)\geq 0.9.$
We henceforth condition on the event $\{\tau\leq H/5,\;s_{\tau}\notin\mathbf{E}_{\text{annotated}}^{\prime}\}$. Applying Lemma 19, we have, with probability at least $0.9$, $\tilde{\pi}$ takes a wrong action at $s_{\tau}$ and transitions to $\mathbf{b^{\prime}}$, and subsequently never takes the expert recovery action at $\mathbf{b^{\prime}}$. Therefore the trajectory remains in $\mathbf{b^{\prime}}$ from time $H/5$ onward, yielding zero reward for at least $\frac{4H}{5}$ steps.
Multiplying all factors, with probability greater than $\frac{1}{2}$,
$J(\pi^{E})-J(\tilde{\pi})\geq\underbrace{0.79}_{\text{reach }\mathbf{E^{\prime}}}\times\underbrace{0.9}_{\text{unannotated }s_{\tau}}\times\underbrace{0.9}_{\text{action errors}}\times\underbrace{0.8H}_{\text{zero reward}}>\frac{H}{2},$
which concludes the proof. ∎
Lemma 17 (Bounded $\mathbf{E^{\prime}}$ coverage).
If the number of expert trajectories satisfies:
$N_{\mathrm{off}}\leq\frac{N_{1}}{160},$
then, with probability at least $\frac{1}{2}$,
$|\mathbf{E}_{\text{annotated}}^{\prime}|\;\leq\;\frac{N_{1}}{10}.$
Proof.
Recall that by Observation 1, $d_{h}^{\pi^{E}}(s)=\frac{\beta}{N_{1}(1+\beta)}=\frac{8}{HN_{1}}$ for $s\in\mathbf{E^{\prime}}$, where $\beta=\frac{8}{H-8}$. Denote $X:=|\mathbf{E^{\prime}}_{\text{annotated}}|$ to be the number of annotated states in $\mathbf{E^{\prime}}$. Consider $N_{\mathrm{off}}$ expert trajectories, each of length $H$, and let $s_{i,h}$ be the state at step $h$ of trajectory $i$. Fix any $s\in\mathbf{E^{\prime}}$. By a union bound over all time indices,
$\Pr(s\in\mathbf{E^{\prime}}_{\text{annotated}})=\Pr\Big(\bigcup_{i=1}^{N_{\mathrm{off}}}\bigcup_{h=1}^{H}\{s_{i,h}=s\}\Big)\leq\sum_{i=1}^{N_{\mathrm{off}}}\sum_{h=1}^{H}\Pr(s_{i,h}=s)=\sum_{i=1}^{N_{\mathrm{off}}}\sum_{h=1}^{H}d_{h}^{\pi^{E}}(s)=\frac{8N_{\mathrm{off}}}{N_{1}}.$
Under the assumption $N_{\mathrm{off}}\leq N_{1}/160$, this gives
$\Pr(s\in\mathbf{E^{\prime}}_{\text{annotated}})\leq\frac{1}{20}.$
Let $Z_{s}$ be the indicator that $s$ is visited by the expert trajectories, by linearity of expectation,
$\mathbb{E}[X]=\sum_{s\in\mathbf{E^{\prime}}}\mathbb{E}[Z_{s}]=\sum_{s\in\mathbf{E^{\prime}}}\Pr(s\in\mathbf{E^{\prime}}_{\text{annotated}})\leq\frac{N_{1}}{20}.$
Applying Markov’s inequality at the threshold $N_{1}/10$,
$\Pr\Big(X>\frac{N_{1}}{10}\Big)\leq\frac{\mathbb{E}[X]}{N_{1}/10}\leq\frac{N_{1}/20}{N_{1}/10}=\frac{1}{2}.$
Equivalently,
$\Pr\Big(X\leq\frac{N_{1}}{10}\Big)\geq\frac{1}{2}.$
Lemma 18 (First $\mathbf{E^{\prime}}$ visit).
For the MDP $\mathcal{M}$, for any policy $\pi$ that agrees with $\pi^{\mathrm{E}}$ on $\mathbf{E}$:
$\Pr_{\mathcal{M},\pi}\big(\exists h\in[H/5],s_{t}\in\mathbf{E}^{\prime}\;\big)\geq 0.79$
Proof.
Since $\pi$ agrees with $\pi^{\mathrm{E}}$ on $\mathbf{E}$,
$\Pr_{\mathcal{M},\pi}\big(\exists h\in[H/5],s_{t}\in\mathbf{E}^{\prime}\;\big)=\Pr_{\mathcal{M},\pi^{\mathrm{E}}}\big(\exists h\in[H/5],s_{t}\in\mathbf{E}^{\prime}\;\big).$
It suffices to consider the expert policy $\pi^{\mathrm{E}}$’s visitation. By Observation 1, the state visitation distribution for $\pi^{\mathrm{E}}$ satisfies that for all $h\in[H]$,
$d^{\pi^{E}}_{h}(s)=\begin{cases}\frac{1}{N_{0}(1+\beta)},\;&s^{\prime}\in\mathbf{E},\\ \frac{\beta}{N_{1}(1+\beta)},\;&s^{\prime}\in\mathbf{E},\\ 0,&\text{otherwise}.\end{cases}$
The probability of no $s\in\mathbf{E^{\prime}}$ visit in $\frac{H}{5}$ steps is:
$\frac{1}{1+\beta}\cdot(1-\beta)^{\frac{H}{5}-1}=$ $\frac{1}{(1+\beta)(1-\beta)}\left(\frac{1}{1+\beta}\right)^{\frac{H}{5}}$ (4)
$=$ $\frac{1}{1-\beta^{2}}\left(1-\frac{8}{H}\right)^{\frac{H}{5}}$
$\leq$ $\frac{1}{1-\frac{64}{(H-8)^{2}}}e^{-\frac{8}{H}\cdot\frac{H}{5}}$
[…]
where we apply $\beta=\frac{8}{H-8}$, $1-x\leq e^{-x}$, and $H\geq 50$. Thus:
$\Pr_{\mathcal{M},\pi^{\mathrm{E}}}\big(\exists h\in[H/5],s_{t}\in\mathbf{E}^{\prime}\;\big)>1-0.2095=0.79$
Lemma 19 (Action Errors on Unannotated States).
Let ${\mathcal{S}}_{\text{unannotated}}\subseteq{\mathcal{S}}$ be the set of unannotated states for a policy $\pi$. Consider MDP $\mathcal{M}$ with action space size $A\geq 10H$.
$\Pr\big(\forall h\in[H]:s_{h}\notin{\mathcal{S}}_{\text{annotated}}\implies\ \pi(s_{h})\neq\pi^{E}(s_{h})\big)\;\geq\;0.9.$
Proof.
Since for any unannotated state $s$, the learner’s policy selects the expert action with probability exactly $\frac{1}{A}$. Condition on any realization with $|H_{\text{unannotated}}|\leq H$. By union bound,
$\Pr\big(\exists h\in[H]:\ s_{h}\in{\mathcal{S}}_{\text{unannotated}}\text{ and }\pi(s_{h})=\pi^{E}(s_{h})\big)\leq\frac{H}{A}\leq\frac{H}{10H}=0.1.$
Thus
$\Pr\big(\forall h\in[H]:s_{h}\notin{\mathcal{S}}_{\text{annotated}}\implies\ \pi(s_{h})\neq\pi^{E}(s_{h})\big)\;\geq\;0.9.$
D.2 Lower Bound for Stagger
Throughout the proof, denote by $\mathbf{E}_{\text{annotated}}$ the final set of states in $\mathbf{E}$ on which Stagger have requested expert annotations.
Lemma 20 (Stagger suboptimality lower bound).
Consider the MDP $\mathcal{M}$ from Section 4.2 with $H\geq 50$, $A\geq 10H$, and $\beta=\frac{8}{H-8}$. If Stagger collects no more than
$N_{\mathrm{int}}\leq\frac{HN_{0}}{12}$
interactive state-wise annotations, then, with probability at least $\frac{1}{2}$, the returned policy $\hat{\pi}$ suffers suboptimality at least
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\geq\frac{H}{2}.$
Proof.
By Lemma 21, if Stagger collects at most $\frac{HN_{0}}{12}$ interactive state-wise annotations as above, then with probability at least $\frac{1}{2}$, $|\mathbf{E}_{\text{annotated}}|$, the number of distinct states in $\mathbf{E}$ annotated is fewer than $N_{0}/3$. Consider a random rollout of $\hat{\pi}$, define the following events:
$F_{1}$ $:=\left\{s_{1}\notin\mathbf{E}_{\text{annotated}}\right\},$
$F_{2}$ $:=\{a_{1}\neq\pi^{\mathrm{E}}(s_{1})\}.$
We now lower bound their probabilities:
$\Pr_{\hat{\pi}}(F_{1})$ $\geq\frac{2}{3(1+\beta)},$
$\Pr_{\hat{\pi}}(F_{2}\mid F_{1})$ $\geq 1-\frac{1}{A}\geq 1-\frac{1}{10H}.$
Conditioned on the two events, the agent will get trapped at state $\mathbf{b}$ from step 2 on, and thus its conditional expected return satisfies
$\mathbb{E}_{\hat{\pi}}\left[\sum_{h=1}^{H}r_{h}\mid F_{1},F_{2}\right]\leq 1.$
Also, by the definition of the reward function $R_{1}$, $J(\pi^{E})=H$. Thus,
[…]
$\geq$ $\mathbb{E}_{\hat{\pi}}\left[H-\sum_{h=1}^{H}r_{h}\mid F_{1},F_{2}\right]P_{\hat{\pi}}(F_{1},F_{2})$
$\geq$ $\frac{2}{3}\cdot\frac{H-8}{H}\cdot\frac{10H-1}{10H}\cdot(H-1)\geq\frac{H}{2}$
Where we use our setting of $\beta=\frac{8}{H-8}$ and apply $H\geq 50$.
Lemma 21 (Bounded $\mathbf{E}$ coverage under Stagger).
Suppose Stagger collects at most
$N_{\mathrm{int}}\leq\frac{HN_{0}}{12}$
interactive annotations. Then,
$\Pr\left(|\mathbf{E}_{\text{annotated}}|\geq\frac{N_{0}}{3}\right)\leq\frac{1}{2}.$
Proof.
In this proof, we say that state $s$ is annotated at iteration $i$, if it has been annotated by the expert before iteration $i$ (excluding iteration $i$). Since each iteration of Stagger samples only one state uniformly from the current episode for annotation, we denote the indicator of expert annotating an unannotated state from $\mathbf{E}$ at iteration $i$ as $X_{i}\in\{0,1\}$. With this notation, we have the total number of annotated states by the end of iteration $N_{\mathrm{int}}$ as
$|\mathbf{E}_{\text{annotated}}|=\sum_{i=1}^{N_{\mathrm{int}}}X_{i}.$
Let $\mathcal{F}_{j}$ be the sigma-algebra generated by all information seen by Stagger up to iteration $j$. We now upper bound the expected value of $X_{i}$ conditioned on $\mathcal{F}_{i-1}$, for each $i$.
Denote by $Y_{i}$ the number of unannotated states in $\mathbf{E}$ visited by round $i$’s rollout (by policy $\pi^{i}$). We claim that conditioned on $\mathcal{F}_{i-1}$, $Y_{i}$ is stochastically dominated by a geometric distribution with parameter $\frac{A-1}{A}$. Indeed, whenever an unannotated state in $\mathbf{E}$ is encountered when rolling out $\pi^{i}$, the probability that the agent takes a wrong action is $\frac{A-1}{A}$; if so, the agent gets absorbed to $\mathbf{b}$ immediately, and thus never sees any new unannotated states in this episode. In summary,
$\mathbb{E}[Y_{i}\mid\mathcal{F}_{i-1}]\leq\mathbb{E}_{Z\sim\mathrm{Geometric}(\frac{A-1}{A})}[Z]=\frac{A}{A-1}\leq 2.$
Since the state sampled for expert annotation is uniformly at random from the trajectory, conditioned on $Y_{i}$, the probability that it lands on a an unannotated state in $\mathbf{E}$ is at most $\frac{Y_{i}}{H}$. Hence,
$\mathbb{E}[X_{i}\mid\mathcal{F}_{i-1}]=\mathbb{E}\left[\mathbb{E}[X_{i}\mid Y_{i},\mathcal{F}_{i-1}]\mid\mathcal{F}_{i-1}\right]=\mathbb{E}\left[\frac{Y_{i}}{H}\mid\mathcal{F}_{i-1}\right]\leq\frac{2}{H}.$
By linearity of expectation:
$\mathbb{E}[|\mathbf{E}_{\text{annotated}}|]=\sum_{i=1}^{N_{\mathrm{int}}}\mathbb{E}[X_{i}]\leq\frac{2N_{\mathrm{int}}}{H}.$
Applying Markov’s inequality:
$\Pr(|\mathbf{E}_{\text{annotated}}|\geq N_{0}/3)\leq\frac{\mathbb{E}[X]}{N_{0}/3}\leq\frac{6N_{\mathrm{int}}}{HN_{0}}\leq\frac{1}{2}.$
where the last inequality is by our assumption that $N_{\mathrm{int}}\leq\frac{HN_{0}}{12}$. This completes the proof. ∎
D.3 Upper Bound for Warm-Stagger
Lemma 22 (Hybrid IL achieves expert performance under $R_{1}$).
Consider the MDP $\mathcal{M}$ and expert policy $\pi^{\mathrm{E}}$ as above, then, with probability at least $1/2$, Warm-Stagger outputs a policy that achieves expert performance using $N_{\mathrm{off}}=O\left(\frac{N_{0}}{H}\log(N_{0})\right)$ offline expert trajectories and $N_{\mathrm{int}}\leq 3$ interactive annotations.
Proof.
We divide the proof into four parts. In this proof, we denote by $\mathbf{E}_{\text{annotated}}$ and $\mathbf{E}_{\text{annotated}}^{\prime}$ the subsets of $\mathbf{E}$ and $\mathbf{E^{\prime}}$, respectively, that are annotated by the $N_{\mathrm{off}}$ offline expert demonstration trajectories.
First
, we state a high probability event for the $N_{\mathrm{off}}$ offline expert demonstration trajectories to provide annotations on all states in $\mathbf{E}$. Define event
$F_{3}:=\{\mathbf{E}_{\text{annotated}}=\mathbf{E}\}.$
By Lemma 23, the choice
$N_{\mathrm{off}}=\frac{N_{0}}{(1-\beta)H}\log\!\Big(10N_{0}\Big)$
ensures that $\Pr(F_{3})\geq 0.9$. On event $F_{3}$, the learner takes the correct action on every state in $\mathbf{E}$.
Next
, we define the event under which only a small fraction of $\mathbf{E^{\prime}}$ is covered by the expert. With $H\geq 50$, $H\geq\frac{5}{4}\log(10N_{0})$, $\beta=\frac{8}{H-8}$, $N_{1}\geq 500$, and $N_{1}\geq 160N_{0}$, the number of offline trajectories satisfies
$N_{\mathrm{off}}=\frac{N_{0}}{(1-\beta)H}\log\!\Big(10N_{0}\Big)\leq\frac{4}{5(1-\beta)}N_{0}=\frac{4(H-8)}{5(H-16)}N_{0}\leq N_{0}<\frac{N_{1}}{160},$
where we apply $H\geq\frac{5}{4}\log(10N_{0})$ and $\frac{H-16}{H-8}<\frac{4}{5}$ when $H\geq 50$. Thus, by Lemma 17,
$\mathbb{E}\!\left[|\mathbf{E}_{\text{annotated}}^{\prime}|\right]\leq\frac{N_{1}}{20}.$
Applying Markov’s inequality at threshold $\frac{N_{1}}{4}-2$ gives
$\Pr\Big(|\mathbf{E}_{\text{annotated}}^{\prime}|>\frac{N_{1}}{4}-2\Big)\leq\Pr\Big(|\mathbf{E}_{\text{annotated}}^{\prime}|>\frac{N_{1}}{4}-\frac{2N_{1}}{500}\Big)\leq\frac{25}{121}<0.21.$
Define event
$F_{4}:=\left\{|\mathbf{E}_{\text{annotated}}^{\prime}|\leq\frac{N_{1}}{4}-2\right\}.$
Then $\Pr(F_{4})\geq 0.79$.
Third
, we show that under events $F_{3}$ and $F_{4}$, the interactive phase of Warm-Stagger annotates $\mathbf{b^{\prime}}$ with probability greater than 0.4 in each of the first three rollouts, such that $\mathbf{b^{\prime}}$ is annotated within three expert annotations with probability greater than 0.78.
Conditioned on $F_{3}$, the learner acts optimally on all states in $\mathbf{E}$. Conditioned on $F_{4}$, at least $\frac{3}{4}$ of the states in $\mathbf{E^{\prime}}$ are unannotated for each of the first three rollouts. Let $\tau$ be the first index with $s_{\tau}\in\mathbf{E^{\prime}}$. By Lemma 18,
$\Pr(\tau\leq H/5)\geq 0.79.$
Since $s_{\tau}$ is drawn uniformly from $\mathbf{E^{\prime}}$ and at most $N_{1}/4$ states are annotated under $B$,
$\Pr(s_{\tau}\notin\mathbf{E}_{\text{annotated}}^{\prime}\mid F_{3},F_{4},\tau\leq H/5)\;\geq\;\frac{3}{4}.$
With $A\geq 10H$, by Lemma 19,
$\Pr\!\left(\pi(s_{\tau})\neq\pi^{E}(s_{\tau})\mid s_{\tau}\notin\mathbf{E}_{\text{annotated}}^{\prime},F_{3},F_{4},\tau\leq H/5\right)\geq 0.9,$
so with probability at least $0.9$ the learner transitions to $\mathbf{b^{\prime}}$ at time $\tau+1$ and stays there till the end of episode, accumulating no less than $0.8H$ $\mathbf{b^{\prime}}$ states with $H\geq 50$. Since Warm-Stagger samples one state uniformly from each rollout for annotation, this gives 0.8 probability of sampling $\mathbf{b^{\prime}}$ in this specific case.
Combining these factors, we obtain
[…]
Define event
$F_{5}:=\{\mathbf{b^{\prime}}\text{ is annotated within 3 rollouts}\}.$
Conditioned on $F_{3}$ and $F_{4}$, the three rollouts has in each rollout the probability of sampling $\mathbf{b^{\prime}}$ is at least $0.4$. Hence
$\Pr(F_{5}\mid F_{3},F_{4})\;\geq\;1-(1-0.4)^{3}=1-0.6^{3}>0.78.$
Combining the three events, we have
$\Pr(F_{3}\cap F_{4}\cap F_{5})\;\geq\;\Pr(F_{3}\cap F_{4})\,\Pr(F_{5}\mid F_{3},F_{4})\;\geq\;(0.9+0.79-1)\times 0.78>0.5.$
Finally
, on $F_{3}\cap F_{4}\cap F_{5}$, all states in $\mathbf{E}$ and the reset state $\mathbf{b^{\prime}}$ are annotated, under reward function $R_{1}$, the learner receives reward 1 at any step if it is in $\mathbf{E}$ or $\mathbf{E^{\prime}}$ or it is in $\mathbf{b^{\prime}}$ and takes the recovery action same as the expert. Since the learner now behaves like $\pi^{\mathrm{E}}$ on all states in $\mathbf{E}$ and can successfully recovers in $\mathbf{B^{\prime}}$, its total return is the same as the expert, which concludes the proof.
Lemma 23 (Coverage of expert with $N_{\text{off}}$ trajectories).
Consider the MDP $\mathcal{M}$. For $N_{\mathrm{off}}$ trajectories of length $H$ rolled out by expert policy $\pi^{\mathrm{E}}$, all $N_{0}$ states in $\mathbf{E}$ are annotated with probability $\geq 1-\delta$ if:
$N_{\mathrm{off}}\;\geq\;\frac{N_{0}}{(1-\beta)H}\log\!\Big(\frac{N_{0}}{\delta}\Big).$
Proof.
Recall that we denote by $\mathbf{E}_{\text{annotated}}$ the set of states in $\mathbf{E}$ visited and annotated by offline expert demonstrations. Fix any $s\in\mathbf{E}$ and concatenate all expert trajectories into a single sequence $\{s_{t}\}_{t=1}^{T}$ of length $T:=HN_{\mathrm{off}}$, ordered from the first state of the first trajectory to the last state of the last trajectory. Let $\mathcal{F}_{t}$ be the sigma-algebra generated by all states $s_{1},\dots,s_{t}$, and define the set of initial indices
$I_{0}:=\{\,1,\;H+1,\;2H+1,\;\dots,\;(N_{\mathrm{off}}-1)H+1\,\}.$
For any $t+1\in I_{0}$, $s_{t+1}$ is drawn from $\rho$, so
$\Pr(s_{t+1}=s\mid\mathcal{F}_{t})=\rho(s)=\frac{1}{(1+\beta)N_{0}}.$
For $t+1\notin I_{0}$, we have $s_{t}\in\mathbf{E}\cup\mathbf{E^{\prime}}$ and under $\pi^{E}$:
$\Pr(s_{t+1}=s\mid s_{t})=\begin{cases}\dfrac{1-\beta}{N_{0}},&s_{t}\in\mathbf{E},\\[2.0pt] \dfrac{1}{N_{0}},&s_{t}\in\mathbf{E^{\prime}}.\end{cases}$
Therefore, in all cases,
$\Pr(s_{t+1}=s\mid\mathcal{F}_{t})\;\geq\;\frac{1-\beta}{N_{0}},\qquad\Pr(s_{t+1}\neq s\mid\mathcal{F}_{t})\;\leq\;1-\frac{1-\beta}{N_{0}}.$
Let $A_{t}:=\{s_{1}\neq s,\dots,s_{t}\neq s\}$. Then
$\Pr(A_{t+1})\leq\Big(1-\frac{1-\beta}{N_{0}}\Big)\Pr(A_{t}),$
so by induction,
$\Pr(A_{T})\leq\Big(1-\frac{1-\beta}{N_{0}}\Big)^{T}\leq\exp\!\Big(-\frac{(1-\beta)T}{N_{0}}\Big)=\exp\!\Big(-\frac{(1-\beta)HN_{\mathrm{off}}}{N_{0}}\Big).$
Therefore,
$\Pr\big(s\notin\mathbf{E}_{\text{annotated}}\big)=\Pr(A_{T})\leq\exp\!\Big(-\frac{(1-\beta)HN_{\mathrm{off}}}{N_{0}}\Big).$
By a union bound over all $s\in\mathbf{E}$,
$\Pr(\exists s\in\mathbf{E},s\notin\mathbf{E}_{\text{annotated}})\leq N_{0}\exp\!\Big(-\frac{(1-\beta)HN_{\mathrm{off}}}{N_{0}}\Big),$
which is at most $\delta$ whenever
$N_{\mathrm{off}}\;\geq\;\frac{N_{0}}{(1-\beta)H}\log\!\Big(\frac{N_{0}}{\delta}\Big).$
Appendix E Additional Guarantees for DAgger Variant Without Recoverability Assumption
In this section, we revisit and conduct a refined analysis of another variant of DAgger with trajectory-wise annotations. We show that without the recoverability assumption, an interactive IL algorithm has sample complexity no worse than that of behavior cloning. This result complements prior work [foster2024behavior] that analyzes a different version of DAgger, which they proved to have a worse sample complexity guarantee than behavior cloning.
E.1 Additional Notations and Useful Distance Measures
In line with [foster2024behavior], we consider another oracle that models interacting with the demonstration expert: the trajectory-wise demonstration oracle $\mathcal{O}^{\mathrm{Traj}}$ that takes into a state-sequence $s_{1:H}$ and returns $a^{*}_{1:H}\sim\pi^{\mathrm{E}}(\cdot\parallel s_{1:H})$. Different from other sections, in the following we adopt Markovian policy, which is a collection of $H$ mappings from states to probability distributions over actions $\pi=\{\pi_{h}:{\mathcal{S}}\rightarrow\Delta(\mathcal{A})\}_{h=1}^{H}$.
We here provide a formal definition of first-step policy mixing—used in the definition of $\hat{\pi}$ (see Algorithm 1, Algorithm 2), which is a common technique (e.g., [sun2017deeply, syed2008apprenticeship, osband2016deep]), defined as follows:
Definition 24 (First-step mixing of $\mathcal{B}$).
$\Pi_{\mathcal{B}}:=\left\{\pi_{u}:u\in\Delta(\mathcal{B})\right\}$, where policy $\pi_{u}$ is executed in an an episode of an MDP $\mathcal{M}$ by: draw $\pi\sim u$ at the beginning of the episode, and execute policy $\pi$ throughout the episode.
Importantly, $\pi_{u}$ is not a stationary policy (note its difference with $\bar{\pi}_{u}$ in Definition 1); as a result, $a_{1:H}$ are dependent conditioned on $s_{1:H}$, while $a_{1:H}$ are only conditionally independent given $s_{1:H}$ and the random policy $\pi$ drawn.
Additionally, we use $\pi(\cdot\parallel s_{1:H})$ to denote the causally-conditioned probability of action sequence $a_{1:H}$ induced by $\pi$, given state sequence $s_{1:H}$ [ziebart2010modeling].  [^2] The use of $\parallel$ highlights its distinction from standard conditioning on $s_{1:H}$. To elaborate:
  • For Markovian policy $\pi$, $\pi(a_{1:H}\parallel s_{1:H}):=\prod_{h=1}^{H}\pi_{h}(a_{h}|s_{h})$.
  • For first-step mixing of Markovian policies $\pi_{u}$, $\pi_{u}(\cdot\parallel s_{1:H}):=\sum_{\pi\in\mathcal{B}}u(\pi)\pi(\cdot\parallel s_{1:H})$.
It is well-known that the trajectory distribution induced by Markovian policies and their first-step mixings $\pi$ can be factorized to the product of $\pi(a_{1:H}\parallel s_{1:H})$ and the causally-conditioned probability of the state sequence given the action sequence (Definition 40 and Lemma 41). When it is clear from context, we use shorthand $\pi(s_{1:H})$ for $\pi(\cdot\parallel s_{1:H})$.
In the following, we present another 2 useful distance measures for pair of policies.
Definition 25 (Trajectory-wise $L_{\infty}$-semi-metric [foster2024behavior]).
For a pair of Markovian policies $\pi$ and $\pi^{\prime}$, define their trajectory-wise $L_{\infty}$-semi-metric as
$\rho(\pi\parallel\pi^{\prime}):=\mathbb{E}^{\pi}\mathbb{E}_{a^{\prime}_{1:H}\sim\pi^{\prime}(s_{1:H})}\left[\mathbb{I}\left\{\exists h:a_{h}\neq a^{\prime}_{h}\right\}\right].$
$\rho(\pi\parallel\pi^{\prime})$ is the probability of any action taken by $\pi^{\prime}$ deviating from actions in trajectories induced by $\pi$, which is symmetric [foster2024behavior]. A bound on $\rho(\pi\parallel\pi^{\mathrm{E}})$ leads to straightforward performance difference guarantee: $J(\pi^{\mathrm{E}})-J(\pi)\leq R\cdot\rho(\pi\parallel\pi^{\mathrm{E}})$ [foster2024behavior] (Lemma 37).
Definition 26 (Decoupled Hellinger distance).
For a pair of Markovian policies $\pi$ and $\pi^{\prime}$, define their decoupled Hellinger distance as $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\prime}(s_{1:H}))\right]$.
Similarly, $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\prime}(s_{1:H}))\right]$ denotes the expected Hellinger distance between the distribution of actions $\pi(s_{1:H})$ and $\pi^{\prime}(s_{1:H})$ on state sequence $s_{1:H}$ visited by $\pi$. This allows decoupled analysis for state and action sequences, which is useful for the proof of Theorem 27.
E.2 Interactive IL Matches Offline IL on Trajectory-wise Annotation
Next, we consider the trajectory-wise sampling model. We present Tragger, another DAgger variant, namely Algorithm 3 and provide its sample complexity bounds.
Algorithm 3 Tragger: DAgger with trajectory-wise annotation oracle
1:Input: MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$ with decision space $\Delta(\mathcal{B})$ and benchmark set $\left\{e_{\pi}:\pi\in\mathcal{B}\right\}$.
2:for $n=1,\dots,N$ do
3:  Query $\mathbb{A}$ and receive $u^{n}\in\Delta(\mathcal{B})$.
4:  Execute $\pi^{n}:=\pi_{u^{n}}$ and sample $s^{n}_{1:H}$ following $\mathbb{P}^{\pi^{n}}$. Query $\mathcal{O}^{\mathrm{Traj}}$ for $a_{1:H}^{*,n}=\pi^{\mathrm{E}}(s^{n}_{1:H})$.
5:  Update $\mathbb{A}$ with loss function
$\ell^{n}(\pi):=\log\left(\frac{1}{\pi^{n}(a_{1:H}^{*,n}\parallel s_{1:H}^{n})}\right).$ (5)
6:end for
7: Output $\hat{\pi}$, the first-step uniform mixture of policies in $\left\{\pi^{1},\ldots,\pi^{n}\right\}$.
Algorithm 3 uses first-step mixing policies $\pi_{u}\in\Pi_{\mathcal{B}}$ (recall Definition 24). At round $n$, it rolls out $\pi^{n}=\pi_{u^{n}}$ whose mixing weight $u^{n}$ is obtained from an online learning oracle $\mathbb{A}$ and samples a full state sequence $s^{n}_{1:H}$. Same as Algorithm 1, Algorithm 3 also requires $\mathbb{A}$ to have decision space $\Delta(\mathcal{B})$ and benchmark set $\left\{e_{\pi}:\pi\in\mathcal{B}\right\}$. It then requests expert’s trajectory-wise annotation $a^{*,n}_{1:H}$ and updates $\mathbb{A}$ by $\ell^{n}(\pi)$ (Eq. (5)). At the end of round $N$, the uniform first-step mixing of $\{\pi^{n}\}_{n=1}^{N}$ is returned, which is equivalent to returning $\pi_{\hat{u}}$, where $\hat{u}:=\frac{1}{H}\sum_{n=1}^{N}u^{n}$. We provide the following performance guarantee of Algorithm 3:
Theorem 27.
If Algorithm 3 is run with a deterministic expert policy $\pi^{E}$, a policy class $\mathcal{B}$ such that realizability holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm, then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq 2R\frac{\log(B)+2\log(1/\delta)}{N}.$
Theorem 27 shows that the interactive IL Algorithm 3 matches the trajectory-wise sample complexity of behavior cloning in [foster2024behavior]. In contrast, prior state-of-the-art analysis of interactive IL algorithms [foster2024behavior, Appendix C.2] gives sample complexity results that are in general worse than behavior cloning.  [^3] For [foster2024behavior, Appendix C.2]’s sample complexity to improve over behavior cloning, we need $\mu H\max_{h\in[H]}\log|\mathcal{B}_{h}|$ to be significantly smaller $R\log|\mathcal{B}|$ (where $\mathcal{B}_{h}$ is the projection of $\mathcal{B}$ onto step $h$). This may require the strong condition that $\mu<R/H<1$ in the practically-popular parameter sharing settings ($|\mathcal{B}_{h}|=|\mathcal{B}|$).
For the proof of Theorem 27, we introduce a new notion of decoupled Hellinger estimation error:
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}:=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}$ decouples the dependence between the state sequence and the distribution of action sequence induced by the learner. Perhaps surprisingly, it is compatible with non-Markovian first-step mixing of policies, while still being well-behaved enough to be translated to a policy suboptimality guarantee, which could be of independent interest.
E.3 Decoupling State and Action Sequences by Decoupled Hellinger Distance
In this section, we demonstrate that similar to $D^{2}_{\text{H}}\left(\mathbb{P}^{\pi}(s_{1:H},a_{1:H}),\mathbb{P}^{\pi^{\mathrm{E}}}(s_{1:H},a_{1:H})\right)$ [foster2024behavior], the decoupled Hellinger distance $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right]$ that decouples states and actions is also proportionally lower bounded by a constant factor of $\rho(\pi\parallel\pi^{\mathrm{E}})$. The following two lemma shows that such relationship holds for both Markovian policies and their first-step mixings.
Lemma 28.
Let $\pi^{\mathrm{E}}$ be a deterministic policy, and let $\pi$ be an Markovian policy. Then we have
$\frac{1}{2}\cdot\rho(\pi\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
Lemma 29.
Let $\pi^{\mathrm{E}}$ be a deterministic policy, and let $\pi_{u}$ be a first-step mixing of Markovian policies. Then we have
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
To prove these two lemmas, we first prove a special case, i.e. Lemma 30 with first-step mixing of deterministic policies. To facilitate the proofs, we introduce the following additional notations:
  • Let $\mathcal{B}^{\mathrm{Det}}$ denote the set of all deterministic, Markovian policies. We will use $\nu,\nu^{\prime}$ to denote members of $\mathcal{B}^{\mathrm{Det}}$ and $\nu_{h}(s)$ to denote the action taken by $\nu$ at $(s,h)$ when it is clear from the context.
  • Let $\mathcal{B}^{\mathrm{E}}(s_{1:h})$ represent the subset of $\mathcal{B}^{\mathrm{Det}}$ that agrees with $\pi^{\mathrm{E}}$ on the state sequence $s_{1:h}$.
  • Define $F(\nu;\nu^{\prime};\pi^{\mathrm{E}}):=\sum_{s_{1:H}}\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right],$ which evaluates the probability that $\nu^{\prime}$ disagrees with $\pi^{\mathrm{E}}$ over the distribution of $H$-step state sequences induced by $\pi$.
Lemma 30.
Let $\pi^{\mathrm{E}}$ be a deterministic Markovian policy, and let $\pi_{u}$ be a first-step mixing of deterministic Markovian policies (elements of $\mathcal{B}^{\mathrm{Det}}$). Then we have that
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
The key idea in the following proof is to lower bound $\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right],$ which reflects the asymmetric roles of the two appearances of $\pi_{u}$’s, using a symmetric formulation via function $F$ (as shown in (9)).
Proof.
Recall the first-step mixing policy in Definition 24, we start by rewriting
$\rho(\pi_{u}\parallel\pi^{\mathrm{E}})=$ $\mathbb{E}^{\pi_{u}}\left[\mathbb{I}\left\{\exists h:a_{h}\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}\right]$ (6)
$=$ $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\sum_{s_{1:H}}\mathbb{P}^{\nu}(s_{1:H},a_{1:H})\mathbb{I}\left\{\exists h:a_{h}\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}$
$=$ $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\rho(\nu\parallel\pi^{\mathrm{E}}),$
which is a weighted combination of $\rho(\nu\parallel\pi^{\mathrm{E}})$ for $\nu\in\mathcal{B}^{\mathrm{Det}}$.
Next, we turn to analyzing $D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))$. Since the deterministic expert induces a delta mass distribution over actions, we apply the elementary fact about the Hellinger distance with delta mass distribution stated in Lemma 35, yielding:
$\frac{1}{2}\parallel\pi_{u}(s_{1:H})-\pi^{\mathrm{E}}(s_{1:H})\parallel_{1}\leq D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H})).$
We recall that $\mathcal{B}^{\mathrm{E}}(s_{1:H})$ denotes the subset of $\mathcal{B}^{\mathrm{Det}}$ that agrees with $\pi^{\mathrm{E}}$ on $s_{1:H}$ and define the total weight assigned by $u$ on it as $u(\mathcal{B}^{\mathrm{E}}(s_{1:H})):=\sum_{\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:H})}u(\nu)$. Then,
$\frac{1}{2}\parallel\pi_{u}(s_{1:H})-\pi^{\mathrm{E}}(s_{1:H})\parallel_{1}=1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H})),$
which implies:
$1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H}))\leq D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H})).$ (7)
Therefore, by taking expectation over $s_{1:H}\sim\mathbb{P}^{\pi_{u}}$ in Eq. (7),
$\sum_{s_{1:H}}\mathbb{P}^{\pi_{u}}(s_{1:H})(1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H})))\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$ (8)
We now examine the expression
$\sum_{s_{1:H}}\mathbb{P}^{\pi_{u}}(s_{1:H})(1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H}))).$ (*)
Since $\pi_{u}$ is a first-step mixing of policies in $\mathcal{B}^{\mathrm{Det}}$ with weight $u$, we have $\mathbb{P}^{\pi_{u}}(s_{1:H})=\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\mathbb{P}^{\nu}(s_{1:H})$. This allows us to rewrite ()(*) using the definition of $F(\nu;\nu^{\prime},\pi^{\mathrm{E}})$ as:
$(*)=$ $\sum_{s_{1:H}}\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\mathbb{P}^{\nu}(s_{1:H})\sum_{\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu^{\prime})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]$ (9)
$=$ $\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})\sum_{s_{1:H}}\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]$
$=$ $\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})F(\nu;\nu^{\prime};\pi^{E})$
$=$ $\frac{1}{2}\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})\left(F(\nu;\nu^{\prime};\pi^{E})+F(\nu;\nu^{\prime};\pi^{E})\right),$
where the first three equalities are by algebra and the definition of $F(\nu;\nu;\pi^{E})$. In the last equality, we use the observation that
$\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})F(\nu;\nu^{\prime};\pi^{E})=\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})F(\nu^{\prime};\nu;\pi^{E}).$
By Lemma 10 (stated below),
$(*)\geq$ $\frac{1}{2}\cdot\frac{1}{2}\cdot\sum_{\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{Det}}}u(\nu)u(\nu^{\prime})\left(\rho(\nu\parallel\pi^{\mathrm{E}})+\rho(\nu^{\prime}\parallel\pi^{\mathrm{E}})\right)$
$=$ $\frac{1}{2}\cdot\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\rho(\nu\parallel\pi^{\mathrm{E}})=\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}}).$
Combining the above two inequalities with Eq (8) we conclude the proof by
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq(*)\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
Lemma 31 (Symmetric Evaluation Lemma).
Given deterministic Markovian policies $\nu$, $\nu^{\prime}$, and $\pi^{\mathrm{E}}$, the following holds
$\frac{1}{2}\cdot\left(\rho(\nu\parallel\pi^{\mathrm{E}})+\rho(\nu^{\prime}\parallel\pi^{\mathrm{E}})\right)\leq F(\nu;\nu^{\prime};\pi^{\mathrm{E}})+F(\nu^{\prime};\nu;\pi^{\mathrm{E}}).$ (10)
Proof.
Recall that
$F(\nu;\nu^{\prime};\pi^{\mathrm{E}})+F(\nu^{\prime};\nu;\pi^{\mathrm{E}})=\sum_{s_{1:H}}\left(\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left[\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]\right).$
Throughout the proof, we say that $\nu$ makes a mistake at step $h$, if $\nu_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})$. Then, we can partition all state sequences $s_{1:H}\in{\mathcal{S}}^{H}$ into 4 subsets, $\mathcal{X}_{i}$, indexed by $i\in\{1,2,3,4\}$:
  1. 1.
    $\mathcal{X}_{1}:=\{s_{1:H}\;|\;\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:H})\}$;
  2. 2.
    $\mathcal{X}_{2}:=\{s_{1:H}\;|\;\exists h,s.t.\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\}$;
  3. 3.
    $\mathcal{X}_{3}:=\{s_{1:H}\;|\;\exists h,s.t.\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\}$;
  4. 4.
    $\mathcal{X}_{4}:=\{s_{1:H}\;|\;\exists h,s.t.\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\}\}$.
In words, the four subsets divide state sequences into cases where: (1) both $\nu,\nu^{\prime}$ agree with the $\pi^{\mathrm{E}}$ throughout, (2)&\&(3) one of $\nu,\nu^{\prime}$ makes its first mistake earlier than the other, and (4) $\nu,\nu^{\prime}$ make their first mistake at the same time. It can now be easily seen that each $s_{1:H}\in{\mathcal{S}}^{H}$ lies in exactly one of such $\mathcal{X}_{i}$, and
$\mathcal{X}_{1}\cup\mathcal{X}_{2}\cup\mathcal{X}_{3}\cup\mathcal{X}_{4}={\mathcal{S}}^{H}.$
To see this, consider $h^{\text{err}}$, the first time step $h$ such that one of $\nu$ and $\nu^{\prime}$ disagree with $\pi^{\mathrm{E}}$. If $h^{\text{err}}$ does not exist, then $s_{1:H}\in\mathcal{X}_{1}$. Otherwise, $s_{1:H}$ lies in one of $\mathcal{X}_{2},\mathcal{X}_{3},\mathcal{X}_{4}$ depending on whether $\nu$ and $\nu^{\prime}$ makes mistakes at step $h^{\text{err}}$.
By definition, subset $\mathcal{X}_{1}$ denotes trajectories $s_{1:H}$ where $\nu,\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:H})$, meaning that
$\sum_{s_{1:H}\in\mathcal{X}_{1}}\left(\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left[\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]\right)=0\;.$
For the other 3 sets, i.e. $\mathcal{X}_{i}$ for $i\in\{2,3,4\}$, we can further divide each set based on the time step where the first error occurs, formally:
$\mathcal{X}_{2}^{h}$= $\{s_{1:H}\;|\;\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\};$ (11)
$\mathcal{X}_{3}^{h}$= $\{s_{1:H}\;|\;\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\};$
$\mathcal{X}_{4}^{h}$= $\{s_{1:H}\;|\;\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:h}),\nu\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1}),\nu^{\prime}\in\mathcal{B}^{\mathrm{E}}(s_{1:h-1})\}.$
By definition, each pair of subsets is disjoint and $\cup_{h\in[H]}\mathcal{X}_{i}^{h}=\mathcal{X}_{i}$, for $i=2,3,4$. Note that the determination of whether $s_{1:H}\in\mathcal{X}_{i}^{h}$ only depends on $s_{1:h}$; therefore, $\mathcal{X}_{i}^{h}$ can be represented as $\tilde{\mathcal{X}}_{i}^{h}\times{\mathcal{S}}^{H-h}$, where
$\tilde{\mathcal{X}}_{i}^{h}:=\{s_{1:h}\;|\;s_{1:H}\in\mathcal{X}_{i}^{h}\}.$
Based on this observation, we have
$\sum_{s_{1:H}\in\mathcal{X}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:H})=\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h},s_{h+1:H}\in{\mathcal{S}}^{H-h}}\mathbb{P}^{\nu}(s_{1:H})=\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:h}).$
Furthermore, since deterministic policies $\nu,\nu^{\prime},\pi^{\mathrm{E}}$ agrees with each other for all $\{s_{1:h-1}|s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}\}$,
$\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:h})=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}P_{0}(\mathbf{E^{\prime}})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu_{h}(s_{h}))$ (12)
$=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}P_{0}(\mathbf{E^{\prime}})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu^{\prime}_{h}(s_{h}))=\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu^{\prime}}(s_{1:h}).$
This implies that
$\sum_{s_{1:H}\in\mathcal{X}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:H})=\sum_{s_{1:H}\in\mathcal{X}_{i}^{h}}\mathbb{P}^{\nu^{\prime}}(s_{1:H}),$
and therefore, summing over all $h\in[H]$,
$\sum_{s_{1:H}\in\mathcal{X}_{i}}\mathbb{P}^{\nu}(s_{1:H})=\sum_{s_{1:H}\in\mathcal{X}_{i}}\mathbb{P}^{\nu^{\prime}}(s_{1:H}).$
Now, for $\mathcal{X}_{2}$, we have
$\sum_{s_{1:H}\in\mathcal{X}_{2}}\left(\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left[\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]\right)\geq\sum_{s_{1:H}\in\mathcal{X}_{2}}\mathbb{P}^{\nu}(s_{1:H})\;,$ (13)
where we apply the fact that for all $s_{1:H}\in\mathcal{X}_{2}$, $\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})$, and dropping the second term which is nonnegative.
Similarly, for $\mathcal{X}_{3}$, we have that
$\sum_{s_{1:H}\in\mathcal{X}_{3}}\left(\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left[\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]\right)\geq\sum_{s_{1:H}\in\mathcal{X}_{3}}\mathbb{P}^{\nu^{\prime}}(s_{1:H})=\sum_{s_{1:H}\in\mathcal{X}_{3}}\mathbb{P}^{\nu}(s_{1:H}).$ (14)
Finally, for $\mathcal{X}_{4}$, we use the fact that for $s_{1:H}\in\mathcal{X}_{4}$, $\nu,\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})$ and obtain
$\sum_{s_{1:H}\in\mathcal{X}_{4}}\left(\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left[\nu^{\prime}\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left[\nu\notin\mathcal{B}^{\mathrm{E}}(s_{1:H})\right]\right)$ (15)
$=$ $\sum_{s_{1:H}\in\mathcal{X}_{4}}(\mathbb{P}^{\nu}(s_{1:H})+\mathbb{P}^{\nu^{\prime}}(s_{1:H}))\geq\sum_{s_{1:H}\in\mathcal{X}_{4}}\mathbb{P}^{\nu}(s_{1:H}).$
Now, we combine Eqs. (13), (14), (15) and observe that
$\sum_{s_{1:H}\in\mathcal{X}_{2}}\mathbb{P}^{\nu}(s_{1:H})+\sum_{s_{1:H}\in\mathcal{X}_{3}}\mathbb{P}^{\nu}(s_{1:H})+\sum_{s_{1:H}\in\mathcal{X}_{4}}\mathbb{P}^{\nu}(s_{1:H})\geq$ $\frac{1}{2}\sum_{s_{1:H}\in\mathcal{X}_{2}\cup\mathcal{X}_{3}\cup\mathcal{X}_{4}}\left(\mathbb{P}^{\nu}(s_{1:H})+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\right),$ (16)
which implies
$F(\nu;\nu^{\prime};\pi^{\mathrm{E}})+F(\nu^{\prime};\nu;\pi^{\mathrm{E}})\geq\frac{1}{2}\sum_{s_{1:H}\in\mathcal{X}_{2}\cup\mathcal{X}_{3}\cup\mathcal{X}_{4}}\left(\mathbb{P}^{\nu}(s_{1:H})+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\right).$ (17)
Based on the definitions of $\mathcal{X}_{2},\mathcal{X}_{3},\mathcal{X}_{4}$ and $\rho(\cdot\parallel\cdot)$,
$\sum_{s_{1:H}\in\mathcal{X}_{2}\cup\mathcal{X}_{3}\cup\mathcal{X}_{4}}\left(\mathbb{P}^{\nu}(s_{1:H})+\mathbb{P}^{\nu^{\prime}}(s_{1:H})\right)=$ $\sum_{s_{1:H}}\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left\{\exists h:\nu_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\;\text{or}\;\nu^{\prime}_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}$ (18)
$+\sum_{s_{1:H}}\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left\{\exists h:\nu_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\;\text{or}\;\nu^{\prime}_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}$
$\geq$ $\sum_{s_{1:H}}\mathbb{P}^{\nu}(s_{1:H})\mathbb{I}\left\{\exists h:\nu_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}$
$+\sum_{s_{1:H}}\mathbb{P}^{\nu^{\prime}}(s_{1:H})\mathbb{I}\left\{\exists h:\nu^{\prime}_{h}(s_{h})\neq\pi^{\mathrm{E}}_{h}(s_{h})\right\}$
$=$ $\rho(\nu\parallel\pi^{\mathrm{E}})+\rho(\nu^{\prime}\parallel\pi^{\mathrm{E}}),$
where $s_{1:H}\in\mathcal{X}_{2}\cup\mathcal{X}_{3}\cup\mathcal{X}_{4}$ implies either $\nu$ or $\nu^{\prime}$ disagrees with $\pi^{\mathrm{E}}$, while the inequality relaxes the condition by splitting it into separate contributions for $\nu$ and $\nu^{\prime}$.
We conclude the proof by plugging (18) into (17). ∎
E.3.1 Proof of Lemma 28
Proof.
The key to this proof is showing that any Markovian policy $\pi$ is equivalent, in terms of action distribution on any state sequence, to a first-step mixing of a set of deterministic Markovian policies. This leads to equivalence on trajectory distribution and decoupled Hellinger distance. To clarify further, we present the following claim, which allows us to apply guarantees for mixtures of deterministic policies in Lemma 30.
Claim 32.
For a Markovian policy $\pi$, there exists a first-step mixing of deterministic policy $\pi_{u}$ such that for any $s_{1:H}\in{\mathcal{S}}^{H}$, 1. $\pi(s_{1:H})=\pi_{u}(s_{1:H})$, and 2. $\mathbb{P}^{\pi}(s_{1:H})=\mathbb{P}^{\pi_{u}}(s_{1:H})$.
Given an MDP with finite state space size $S$ and action space size $A$, the set of all deterministic, Markovian policies, denoted by $\mathcal{B}^{\mathrm{Det}}$, contains $A^{SH}$ deterministic policies, which can be indexed by a tuple of actions $(a_{h,s})_{h\in[H],\,s\in\mathcal{S}}$.
To construct policy $\pi_{u}$, we will set the weight vector $u$ such that its weight on policy $\nu$ indexed by $(a_{h,s})_{h\in[H],s\in{\mathcal{S}}}$ as:
$u(\nu)=\prod_{h=1}^{H}\prod_{s\in{\mathcal{S}}}\pi_{h}(a_{h,s}|s)$ (19)
It can be easily verified by that $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)=1$.
We now verify the first item. By first-step mixing, we rewrite $\pi_{u}(a_{1:H}\parallel s_{1:H})$ as
$\pi_{u}(a_{1:H}\parallel s_{1:H})=$ $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\prod_{h=1}^{H}\nu(a_{h}|s_{h})$ (20)
$=$ $\sum_{(a^{\prime}_{h,s})_{h\in[H],s\in{\mathcal{S}}}}\prod_{h=1}^{H}\prod_{s\in{\mathcal{S}}}\pi_{h}(a^{\prime}_{h,s}|s)\prod_{h=1}^{H}\mathbb{I}\left[a^{\prime}_{h,s_{h}}=a_{h}\right]$
$=$ $\sum_{(a^{\prime}_{h,s})_{h\in[H],s\neq s_{h}}}\prod_{h=1}^{H}\prod_{s\neq s_{h}}\pi_{h}(a^{\prime}_{h,s}|s)\sum_{(a^{\prime}_{h,s})_{h\in[H],s=s_{h}}}\prod_{h=1}^{H}\pi_{h}(a^{\prime}_{h,s_{h}}|s_{h})\prod_{h=1}^{H}\mathbb{I}\left[a^{\prime}_{h,s_{h}}=a_{h}\right]$
$=$ $\sum_{(a^{\prime}_{h,s})_{h\in[H],s\neq s_{h}}}\prod_{h=1}^{H}\prod_{s\neq s_{h}}\pi_{h}(a^{\prime}_{h,s}|s)\prod_{h=1}^{H}\pi_{h}(a_{h}|s_{h})$
$=$ $\prod_{h=1}^{H}\pi_{h}(a_{h}|s_{h})=\pi(a_{1:H}\parallel s_{1:H}).$
Since this holds for any action sequence $a_{1:H}\in\mathcal{A}^{H}$, we derive the first part of Claim 32 that $\pi(s_{1:H})=\pi_{u}(s_{1:H})$. The second item follows from the first item in combination with Lemma 22.
We conclude that for the $\pi_{u}$ in the statement of the claim,
$\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right]=\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
Finally, the proof follows by applying Lemma 30 to $\pi_{u}$. ∎
E.3.2 Proof of Lemma 29
Proof.
By Claim 32, any Markovian policy can be viewed as a first-step mixing of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$, then any first-step mixing of Markovian policies $\pi_{u}$ can also be viewed as a first-step mixing of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$. The proof follows by applying Lemma 30. ∎
E.4 New Guarantees for DAgger Variant with Trajectory-wise Annotation
Recall that we have defined decoupled Hellinger estimation error:
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
In the following, we first demonstrate that the performance difference between expert and the the uniform first-step mixing of any Markovian policy sequence $\{\pi^{n}\}_{n=1}^{N}$ is upper-bounded by $2R\operatorname{OnEst}^{\mathrm{Traj}}_{N}/N$, and then show the trajectory-wise sample complexity of Algorithm 3 in Theorem 27.
Lemma 33.
For any MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, and sequence of policies $\left\{\pi^{n}\right\}_{n=1}^{N}$, each of which can be Markovian or a first-step mixing of Markovian policies, their first step uniform mixture policy $\hat{\pi}$ satisfies.
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq 2R\cdot\frac{\operatorname{OnEst}^{\mathrm{Traj}}_{N}}{N}.$
Proof.
By Lemma 28 and Lemma 29, for each $\pi^{n}$, whether it is Markovian or a first-step mixing of Markovian policies, the following holds:
$\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right]\geq\frac{1}{2}\rho(\pi^{n}\parallel\pi^{\mathrm{E}}).$
Then, by the definition of $\operatorname{OnEst}^{\mathrm{Traj}}_{N}$,
$\frac{\operatorname{OnEst}^{\mathrm{Traj}}_{N}}{N}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s^{n}_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s^{n}_{1:H}))\right]\geq\frac{1}{2N}\sum_{n=1}^{N}\rho(\pi^{n}\parallel\pi^{\mathrm{E}})=\frac{1}{2}\rho(\hat{\pi}\parallel\pi^{\mathrm{E}}),$
where we apply the fact that $\hat{\pi}$ is a first-step mixing of $\{\pi^{n}\}_{n=1}^{N}$ with uniform weights. Finally, we conclude the proof by applying Lemma 37. ∎
Theorem 34 (Theorem 27 Restated).
If Algorithm 3 is run with a deterministic expert policy $\pi^{E}$, a policy class $\mathcal{B}$ such that realizability holds, and the online learning oracle $\mathbb{A}$ set to exponential weight algorithm (see Proposition 38). Then, with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}\leq\log(B)+2\log(1/\delta),$
and furthermore, the returned $\hat{\pi}$ satisfies
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq 2R\frac{\log(B)+2\log(1/\delta)}{N}.$
Proof.
The proof closely follows Proposition C.2 in [foster2024behavior], tailored for another DAgger variant. However, in this case, we leverage the distribution of the state sequence $s_{1:H}$ instead of the per-step state distribution.
Observe that the log loss functions passed through online learning oracle $\mathbb{A}$, $\ell^{n}(\pi)$ is of the form
$\ell^{n}(\pi)=\log(1/\pi_{u}(a^{n,*}_{1:H}\parallel s^{n}_{1:H}))=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a^{n,*}_{1:H}\parallel s^{n}_{1:H})}\right).$
It can be observed that $\ell^{n}$’s are 1-exp-concave. Therefore, implementing $\mathbb{A}$ using the exponential weights algorithm (Proposition 38) ensures that the following bound holds almost surely:
$\sum_{n=1}^{N}\log(1/\pi^{n}(a^{*,n}_{1:H}\parallel s^{n}_{1:H}))\leq\sum_{n=1}^{N}\log(1/\pi^{\mathrm{E}}(a^{*,n}_{1:H}\parallel s^{n}_{1:H}))+\log(B)=\log(B).$
Then, Lemma 39 with $x^{n}=s^{n}_{1:H}$, $y^{n}=a^{*,n}_{1:H}$, $g_{*}=\pi^{\mathrm{E}}$, and $\mathcal{H}^{n}=\{o^{n^{\prime}}\}_{n^{\prime}=1}^{n}$, where $o^{n}=\left(\mathbf{E^{\prime}}^{n},a_{1}^{n},a_{1}^{*,n},\dots,s_{H}^{n},a_{H}^{n},a_{H}^{*,n}\right)$, implies that with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s^{n}_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s^{n}_{1:H}))\right]\leq\log(B)+2\log(1/\delta).$
Finally, the second part of the theorem follows by applying Lemma 33. ∎
Appendix F Auxiliary Results
Lemma 35.
If $p,q$ are two distributions over some discrete domain $\mathcal{Z}$, and $q$ is a delta mass on an element in $\mathcal{Z}$. Then
$\frac{1}{2}\parallel p-q\parallel_{1}\leq D_{H}^{2}(p\parallel q)\leq\parallel p-q\parallel_{1}$
Lemma 36 (Performance Difference Lemma [Kakade2002ApproximatelyOA][ross2010efficient]).
For two Markovian policies $\pi$ and $\pi^{\mathrm{E}}$ ${\mathcal{S}}\to\Delta(\mathcal{A})$, we have
$J(\pi^{\mathrm{E}})-J(\pi)=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}A_{h}^{E}(s_{h},a_{h})\right],$
where $A_{h}^{E}(s_{h},a_{h}):=Q_{h}^{\pi^{\mathrm{E}}}(s_{h},a_{h})-V_{h}^{\pi^{\mathrm{E}}}(s_{h})$. Furthermore:
  • It holds that (recall Definition 8)
    $J(\pi)-J(\pi^{E})\leq H\cdot\lambda(\pi^{\mathrm{E}}\parallel\pi).$
  • suppose $(\operatorname{\mathcal{M}},\pi^{\mathrm{E}})$ is $\mu$-recoverable, then
    $J(\pi)-J(\pi^{E})\leq\mu\cdot\lambda(\pi\parallel\pi^{\mathrm{E}}).$
Lemma 37 (Lemma D.2. of  [foster2024behavior]).
For all (potentially stochastic) policies $\pi$ and $\pi^{\prime}$, it holds that
$J(\pi)-J(\pi^{\prime})\leq R\cdot\rho(\pi\parallel\pi^{\prime}).$
Proposition 38 (Proposition 3.1 of [cesa2006prediction]).
Suppose $\{\ell^{n}(u)\}_{n=1}^{N}$ is a sequence of $\eta$-exp-concave functions from $\Delta(\mathcal{X})$ to $\mathbb{R}$. For all $x\in\mathcal{X}$, define the weights $w_{x}^{n-1}$ and probabilities $u^{n}(x)$ as follows:
$w_{x}^{n-1}=e^{-\eta\sum_{i=1}^{n-1}\ell_{i}(e_{x})},\quad u^{n}(x)=\frac{w_{x}^{n-1}}{\sum_{x^{\prime}\in\mathcal{X}}w_{x^{\prime}}^{n-1}},$
where $e_{x}$ is the $x$-th standard basis vector in $\mathbb{R}^{|\mathcal{X}|}$. Then, choosing $u^{n}=\{u^{n}(x)\}_{x\in\mathcal{X}}$ (exponential weights used with learning rate $\eta$) satisfies:
$\sum_{n=1}^{N}\ell^{n}(u^{n})\leq\min_{x\in\mathcal{X}}\sum_{n=1}^{N}\ell^{n}(e_{x})+\frac{\log|\mathcal{X}|}{\eta}.$
Lemma 39 (Restatement of Lemma A.14 in [foster2021statistical]).
Under the realizbility assumption, where there exists $g_{\star}:=g_{i_{\star}}\in\mathcal{G}$ such that for all $n\in[N]$,
$y^{n}\sim g_{\star}^{n}(\cdot\mid x^{n})\mid x^{n},\mathcal{H}^{n-1},$
where $\mathcal{H}^{n-1}$ denotes all histories at the beginning of round $n$.
Then, for any estimation algorithm and any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sum_{n=1}^{N}\mathbb{E}_{n-1}\left[D_{\mathcal{H}}^{2}\left(\hat{g}^{n}(x^{n}),g_{\star}^{n}(x^{n})\right)\right]\leq\sum_{n=1}^{N}\left(\ell_{\log}^{n}(\hat{g}^{n})-\ell_{\log}^{n}(g^{n}_{\star})\right)+2\log(\delta^{-1}).$
where $\ell_{\log}^{n}\left(g\right):=\log\left(1/g\left(y^{n}\mid x^{n}\right)\right)$, and $\mathbb{E}_{n}[\cdot]:=\mathbb{E}[\cdot\mid\mathcal{H}^{n}]$.
We have the following well-known lemma for causally-conditioned probability (e.g. [ziebart2010modeling]).
Definition 40.
The causally-conditioned probability of state sequence $s_{1:H}$ given action sequence $a_{1:H-1}$, is defined as
$\mathbb{P}^{\mathcal{M}}(s_{1:H}\parallel a_{1:H-1})=P_{0}(\mathbf{E^{\prime}})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})$
Lemma 41.
For any Markovian policy $\pi$,
$\mathbb{P}^{\pi}(s_{1:H},a_{1:H})=\mathbb{P}^{\mathcal{M}}(s_{1:H}\parallel a_{1:H-1})\cdot\pi(a_{1:H}\parallel s_{1:H}),$ (21)
and for any first-step mixing of Markovian policy $\pi_{u}$,
$\mathbb{P}^{\pi_{u}}(s_{1:H},a_{1:H})=\mathbb{P}^{\mathcal{M}}(s_{1:H}\parallel a_{1:H-1})\cdot{\pi_{u}}(a_{1:H}\parallel s_{1:H}).$ (22)
Proof.
Eq. (21) follows by noticing that both sides are equal to
$P_{0}(\mathbf{E^{\prime}})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\pi_{h}(a_{h}\mid s_{h}).$
Eq. (22) follows by noticing that both sides are equal to
$\sum_{\nu}u(\nu)P_{0}(\mathbf{E^{\prime}})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\nu_{h}(a_{h}\mid s_{h}).$
Appendix G Experiment Details
We compare Warm-Stagger with Behavior Cloning (BC) and Stagger on continuous-control tasks from OpenAI Gym MuJoCo [todorov2012mujoco, brockman2016openai] with episode length $H=1000$.
Infrastructure and Implementation.
All experiments were conducted on a Linux workstation equipped with an Intel Core i9 CPU (3.3GHz) and four NVIDIA GeForce RTX 2080 Ti GPUs. Our implementation builds on the publicly available DRIL framework [brantley2019disagreement] (https://github.com/xkianteb/dril), with modifications to support online learning. The continuous control environments used in our experiments are: “HalfCheetahBulletEnv-v0”, “AntBulletEnv-v0”, “Walker2DBulletEnv-v0”, and “HopperBulletEnv-v0”. We include an anonymous link to our implementation here: https://github.com/liyichen1998/Interactive-and-Hybrid-Imitation-Learning-Provably-Beating-Behavior-Cloning.
Environments and Expert Policies.
We use four MuJoCo environments: Ant, Hopper, HalfCheetah, and Walker2D. The expert policy is a deterministic MLP pretrained via TRPO [schulman2015trust, schulman2017proximal], with two hidden layers of size 64.
Learner Architecture.
The learner uses the same MLP architecture as the expert. Following [foster2024behavior], we use a diagonal Gaussian policy:
$\pi(a\mid s)=\mathcal{N}\left(f_{\theta}(s),\operatorname{diag}(\sigma^{2})\right),$
where $f_{\theta}(s)\in\mathbb{R}^{d_{\mathcal{A}}}$ is the learned mean, and $\sigma\in\mathbb{R}^{d_{\mathcal{A}}}$ is a learnable log-standard deviation vector.
Each model is trained from random initialization using a batch size of 100, a learning rate of $10^{-3}$, and up to 2000 passes over the dataset, with early stopping evaluated every 250 passes using a 20% held-out validation set.
Learning Protocols.
To evaluate the performance of BC against the number of states annotated, we reveal expert state-action pairs sequentially along expert trajectories until the annotation budget is reached. For Stagger, each round it rolls out the latest policy, samples a state uniformly from the trajectory, queries it for the expert action, and updates immediately.
For Warm-Stagger, we begin with BC and switch to Stagger after a predefined number of offline state-action pairs has been used, denoted as $N$. We set $N$ to be 100, 200, or 400 for easier tasks (e.g., Hopper, Ant) and 200, 400, or 800 for harder tasks (e.g., HalfCheetah, Walker2D).
Cost Model and Evaluation.
We assign a cost of 1 to each offline state-action pair and a cost of $C=1$ or $2$ to each interactive query. We run each method for 10 random seeds. For every 50 new state-action pairs collected, we evaluate the current policy by running 25 full-episode rollouts and reporting the average return.
Though the nonrealizable setting is beyond the scope of this work, we expect that some variant of our algorithm can still give reasonable performance, provided that the policy class is expressive enough (so that the approximation error is nonzero but small). For example, [li2023agnostic] observed that with nonrealizable stochastic experts, DAgger variants outperform BC, and exhibit learning curves similar to ours.
G.1 Additional Experiment Plots
We present extended experiment results with larger cost budgets. As shown in Figure 4, we allocate a total annotation budget of 2000 for Hopper and Ant, and 4000 for HalfCheetah and Walker. This complements Figure 3 in the main paper by showing the full training curves without zooming into the stage with small cost budget. The trends are consistent with our earlier observations: Warm-Stagger achieves similar or better sample efficiency compared to Stagger when $C=1$, and clearly outperforms both baselines under the cost-aware setting where $C=2$.
Figure 4: Sample and cost efficiency on MuJoCo tasks. The top row shows expected return vs. number of annotations ($C=1$); the bottom row shows performance under a cost-aware setting ($C=2$). Warm-Stagger (WS) is initialized with 1/20, 1/10, or 1/5 of the samples as offline demonstrations. It matches Stagger in sample efficiency and outperforms the baselines when $C=2$, especially WS(1/5).
G.2 Experiment with MSE Loss
We additionally evaluate our algorithms using mean squared error (MSE) as the loss function for optimization. All training settings remain identical to the main experiments with log loss, except that we use a learning rate of $2.5\times 10^{-4}$. As shown in Figure 5, we observe qualitatively similar results to those under log loss shown in 3, consistent with prior observations in [foster2024behavior], with the added benefit of more stable training dynamics.
Figure 5: Performance comparison under MSE loss across MuJoCo tasks. Results show that Warm-Stagger (WS) achieves comparable sample efficiency and performance to the log loss setting, with improved training stability. Each curve represents the average over 10 seeds.
G.3 Additional Experiments with Algorithm 3
For completeness, we evaluate Tragger and its warm-start variant (Warm-Tragger), as shown in Algorithm 4, on continuous control tasks and the same MDP setup as in Figure 2. The key distinction between Warm-Tragger and Warm-Stagger lies in the annotation mode: the former employs trajectory-wise oracle feedback instead of state-wise annotation, leading to notably different behaviors, as shown in Figure 6.
In particular, for Ant and HalfCheetah, the sample efficiency ($C=1$) of Tragger and Warm-Tragger is significantly worse than that of Stagger due to the cold-start problem: early DAgger rollouts have poor state coverage but must still proceed until the end of each trajectory. In contrast, Stagger samples only a single state per trajectory, thereby better leveraging interactive feedback.
For Hopper and Walker, however, Tragger and Warm-Tragger achieve performance closer to Stagger. These environments feature hard resets when the agent fails (unlikely in Ant and never in HalfCheetah), which truncate poor trajectories and consequently improve sample efficiency.
Overall, these observations suggest a natural middle ground between full-trajectory and single-state annotation—namely, batch queries (e.g., sampling 50 states per trajectory), as explored by [li2023agnostic] with comparable results.
A head-to-head comparison between Tragger and Stagger, as well as between Warm-Tragger and Warm-Stagger, is shown in Figure 7, highlighting the advantage of state-wise over trajectory-wise annotation.
However, this advantage does not hold in general: in the toy MDP in Figure 2, Tragger and Warm-Tragger achieve performance nearly identical to Stagger and Warm-Stagger, as shown in Figure 8.
Algorithm 4 Warm-Tragger: Warm-start Tragger with offline demonstrations
1:Input: MDP $\mathcal{M}$, trajectory-wise expert annotation oracle $\mathcal{O}^{\mathrm{Traj}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$, offline expert dataset $D_{\text{off}}$ of size $N_{\mathrm{off}}$, online budget $N_{\mathrm{int}}$
2: Initialize $\mathbb{A}$ with policy class $\mathcal{B}_{\text{bc}}:=\{\pi\in\mathcal{B}:\pi(s)=\pi^{\mathrm{E}}(s)\ ,\forall s\in D_{\text{off}}\}$.
3:for $n=1,\dots,N_{\mathrm{int}}/H$ do
4:  Query $\mathbb{A}$ and receive $\pi^{n}$.
5:  Execute $\pi^{n}$ and sample $s^{n}_{1:H}$ following $\mathbb{P}^{\pi^{n}}$. Query $\mathcal{O}^{\mathrm{Traj}}$ for $a_{1:H}^{*,n}=\pi^{\mathrm{E}}(s^{n}_{1:H})$.
6:  Update $\mathbb{A}$ with loss function
$\ell^{n}(\pi):=\log\left(\frac{1}{\pi(a_{1:H}^{*,n}\parallel s_{1:H}^{n})}\right).$ (23)
7:end for
8:Output: $\hat{\pi}$, a first-step uniform mixture of $\{\pi^{1},\dots,\pi^{N}\}$.
Figure 6: Sample efficiency of algorithms on MuJoCo tasks, showing expected return vs. number of annotations ($C=1$). Warm-Tragger (WT) is initialized with 1/20, 1/10, or 1/5 of the total annotation budget as offline demonstrations. Specifically, WT($n$) refers to WT with offline demonstrations of total length $n$. Although the performance of WT improves with more offline demonstrations, both Tragger and Warm-Tragger remain inferior to Stagger and, in many cases, even underperform Behavior Cloning, confirming the advantage of state-wise over trajectory-wise annotations.
Figure 7: Head-to-head sample efficiency comparison between Tragger and Stagger, and between Warm-Tragger and Warm-Stagger under equal (since we are talking about comparison here) offline demonstration budgets. Stagger and Warm-Stagger consistently outperform Tragger and Warm-Tragger. The performance gap narrows as the offline budget increases, effectively alleviating the cold-start problem suffered by Tragger.
Figure 8: Similar to Figure 2, we evaluate Tragger and Warm-Tragger (WT) with 200, 800, 3200 offline (state, expert action) pairs in the toy MDP therein. All methods are evaluated under equal total annotation cost with $C=1$. With 800 offline (state, expert action) pairs, WT significantly improves the sample efficiency over the baselines and explores $\mathbf{E^{\prime}}$ more effectively. The performance of Tragger and Warm-Tragger is almost the same as Stagger and Warm-Stagger in Figure 2.